### v1

# Learning to Partially Defer for Sequences

Abstract
In the Learning to Defer (L2D) framework, a prediction model can either make a prediction or defer it to an expert, as determined by a rejector. Current L2D methods train the rejector to decide whether to reject the entire prediction, which is not desirable when the model predicts long sequences. We present an L2D setting for sequence outputs where the system can defer specific outputs of the whole model prediction to an expert in an effort to interleave the expert and machine throughout the prediction. We propose two types of model-based post-hoc rejectors for pre-trained predictors: a token-level rejector, which defers specific token predictions to experts with next token prediction capabilities, and a one-time rejector for experts without such abilities, which defers the remaining sequence from a specific point onward. In the experiments, we also empirically demonstrate that such granular deferrals achieve better cost-accuracy tradeoffs than whole deferrals on Traveling salesman solvers and News summarization models.
Keywords learning to defer, sequences, trustworthy machine learning
1 Introduction
Trustworthiness of AI systems is under increased scrutiny with the increased deployment of such systems in highly consequential areas like healthcare [1, 2, 3] and criminal justice [4, 5, 6]. Hybrid intelligent systems [7, 8, 9, 10] approach this problem through collaborations between either humans and machines or two machines to improve confidence in the system. Learning to Defer (L2D) [11] is one such framework that accommodates these teams by allowing the model to defer to an expert when it is uncertain about a prediction task with the objective of maximizing overall system accuracy while minimizing deferral costs.
Existing L2D methods focus on multiclass classification [12, 13, 14, 15, 16] and scalar regression [17, 18]. However, these methods only support complete deferral of prediction outputs, resulting in costs that scale with the output size, when applied to large sequence outputs [19]. This is especially inefficient if only parts of the model’s prediction are inaccurate. The prevalence of sequences in areas such as drug discovery [20, 21, 22, 23] and language modeling [24, 25, 26] calls for a cost-effective extension of the L2D framework to handle sequential outputs.
To this end, we propose a fine-grained approach to deferral where the system has the flexibility to defer portions of the predicted sequence. For example, in part of speech tagging, this mechanism would only require the expert to tag a few uncertain sentences in a paragraph rather than the whole paragraph, reducing the time cost which increases with the number of words to be tagged. This can, therefore, achieve a more favorable cost-accuracy tradeoff, which can further improve as the deferrals become more precise - for example, at the word-level. Additionally, parts of the prediction task completed by the predictor can also function as contextual clues for the language expert to fill in the deferred parts, further minimizing deferral costs.
While partial deferral is therefore desirable, the granularity of such deferrals depends on the size of the sequence segments the expert can predict at a time. For instance, if the expert is a next-token predictor, it can complete deferred parts of the prediction, as small as a single token. This form of partial deferral that we call token-level deferral can become especially useful when the model’s predictive quality across all parts of the output are dependent. Specifically, in autoregressive prediction, the uncertainty at earlier prediction points can propagate to later points and, in turn, affect the predictive accuracy of the complete sequence [27, 28, 29]. Thus, to achieve a better cost-accuracy tradeoff, the token-level rejector can orchestrate a dynamic collaboration between the predictor and the expert where the predictor receives an expert prediction on an early substructure to predict the next token before the model veers “off-course”.
Although the granularity offered from token-level deferral can improve predictive quality with low deferral costs, this is only feasible with experts that can make predictions one token at a time. That said, partial deferrals can still be implemented as long as the expert is at least capable of completing the predicted sequence given a partial subsequence from the model. We refer to this type of deferral as one-time deferral. For example, Gurobi solvers for Traveling salesman problems lack the ability to iteratively select the next city in the tour, but they can use partial subtours generated by models as constraints to generate the entire tour at a reduced cost. In this case, the one-time rejector would select a point in the sequence from which the expert completes the sequence.
In this paper, we make the following contributions:
  • We present a novel learning to defer framework where the rejector interweaves the expert’s and machine’s predictions to improve cost-accuracy tradeoffs. We divide this setting into token-level and one-time deferrals based on the granularity of the deferrals, thereby accommodating various types of experts.
  • We propose convex surrogate loss functions that offer Bayes consistency guarantees leading to implementable model-based post hoc rejectors for pre-trained predictors for each type of deferral. We further analyze the surrogates theoretically through generalization bounds.
  • We empirically demonstrate improved cost-accuracy tradeoffs when using model-based rejectors for either deferrals against whole sequence rejectors with news summarization and traveling salesman problem solvers.
2 Background
2.1 Learning to Defer
Deferrals, also known as rejections or abstentions, are classified into two types - confidence-based [30, 31, 32, 33] and model-based [34, 11, 12, 14, 13, 15, 17, 18].
Confidence-based methods involve learning a predictor where there is some prescribed cost for rejection; if the confidence score outputted by the learned predictor falls below a cost-dependent threshold, then it is rejected. However, Madras et al. [11] and Cortes et al. [34] characterized the suboptimality of these methods and advocated for the rejector to be a model trained either simultaneously with the predictor or post-hoc: this is the model-based approach. This system seeks to minimize the following loss function:
$\begin{split}{\cal L}_{\text{L2D}}&\left(h,r,e,x,y\right)\\ &=l(y,h(x))\mathbbm{1}_{r(x)<0}+c(x,y,e)\mathbbm{1}_{r(x)\geq 0}\end{split}$ (1)
where $h$ is the predictor, $r$ is the rejector, $e$ is the expert that the model defers to, $l(y,h(x))$ is the model loss, and $c(x,y,e)$ is the cost incurred when deferring to an expert.
While our settings assume access to only one expert, we note that one-time deferral closely resembles L2D with access to predictions from multiple experts [35, 18]. The details of this similarity and its implications are discussed in Section 4.1. In particular, L2D for $E$ experts aims to minimize the following loss function:
$\begin{split}&{\cal L}_{\text{L2DMulti}}\left(h,r,x,y,\{e_{j}\}_{j=1}^{E}\right)\\ &=l(y,h(x))\mathbbm{1}_{r(x)=0}+\sum_{j=1}^{E}c_{j}(x,y,e_{j})\mathbbm{1}_{r(x)=j}\end{split}$ (2)
where $c_{j}(x,y,e_{j})$ is the cost of querying the $j^{\text{th}}$ expert, $e_{j}$. The rejector in this scenario now must decide whether to defer and which expert to defer to.
2.2 Consistent Losses
The direct minimization of Equations 1 and 2 can be challenging due to the non-convex and discontinuous nature of the functions. This can be addressed by optimizing a convex surrogate loss function which, upon minimization, will lead to the same outcome as minimizing the original loss. This property is formally called Bayes Consistency.
Definition 2.1 (Consistency).
Let the risk of a hypothesis $f$ with respect to a loss $l$ be ${\cal R}_{l}(f)=\mathbb{E}_{(x,y)\sim{\cal P}_{{\cal X},{\cal Y}}}[l(f(x),y)]$. Let the optimal risk with respect to $l$ out of all hypotheses in ${\cal F}$ be ${\cal R}_{l}^{*}({\cal F})=\inf_{f\in{\cal F}}{\cal R}_{l}(f)$. A surrogate loss function, $\phi(\cdot)$, is said to be ${\cal F}$-consistent with respect to $l$ if for any $f_{n}\in{\cal F}$, the following property holds true:
$\lim_{n\to\infty}{\cal R}_{\phi}(f_{n})-{\cal R}_{\phi}^{*}({\cal F})=0\Rightarrow\lim_{n\to\infty}{\cal R}_{l}(f_{n})-{\cal R}_{l}^{*}({\cal F})=0$
When ${\cal F}$ is the class of all measurable functions, Bayes consistency is achieved if the excess risk of $l$ can be related to the excess risk of $\phi$ through a non-decreasing function $\Gamma:\mathbb{R}\rightarrow\mathbb{R}$ for which $\Gamma$ is continuous at zero and $\Gamma(0)=0$ in the following manner:
${\cal R}_{l}(f)-{\cal R}_{l}^{*}\leq\Gamma\left({\cal R}_{\phi}(f)-{\cal R}_{\phi}^{*}\right)$
where ${\cal R}_{l}^{*}=\inf_{f}{\cal R}_{l}(f)$ is the optimal risk over the class of all measurable functions. In this paper, we propose loss functions similar to Equations 1 and 2 for both token-level and one-time deferral and then, we show that the above inequality holds true for our surrogate designs for both forms of deferral, proving Bayes consistency.
3 Token-level partial deferral
Let ${\cal Y}\subseteq{\cal V}^{L}$ be the output sequence space where each sequence can have a maximum length $L$ and ${\cal V}$ is the label set that any individual component or token $y_{j}\in y\in{\cal Y}$ of a sequence can belong to. ${\cal V}$ can be a finite vocabulary of tokens in the case of text sequences and be $\mathbb{R}$ when the output is a scalar time series. Let ${\cal X}$ be the feature space. In sequence generation, ${\cal X}\subseteq\mathbb{R}^{d}$ would be a $d$-dimensional feature space, whereas ${\cal X}={\cal V}^{\prime L}$, perhaps, is characterized by a distinct ${\cal V}^{\prime}$ vocabulary set in sequence-to-sequence learning tasks. We denote ${\cal P}_{X,Y}$ to be the data-generating distribution over ${\cal X}\times{\cal Y}$.
Suppose we have a predictor $h$ and an expert $e$ with stronger predictive capabilities than $h$. $e$ takes in $x\in{\cal X}$ and the leftward context $\widehat{y}_{<j}=(\widehat{y}_{1},\cdots,\widehat{y}_{j-1})\in{\cal V}^{j-1}$ as inputs to predict the $j^{\text{th}}$ label in the sequence; repeated calls to $e$ can auto-regressively complete the entire sequence prediction task. $e$ can, for instance, be a human or a large autoregressive model. However, querying $e$ is more expensive than calling $h$. As a result, instead of deferring an entire sequence prediction to the expert, we wish to design an L2D system where experts can complete uncertain tokens in the model’s prediction, facilitating more fine-grained deferrals - referred to as token-level deferrals - in order to achieve a better cost-quality tradeoff.
If $h$ can also perform autoregressive prediction, token-level deferrals can additionally enable the predictor to leverage the predicted token from the expert to predict the next token, fostering a more seamless and adaptive collaboration between the expert and the predictor. This can further improve cost-effectiveness by curbing uncertainty propagation for the predictor, thereby reducing the number of expert calls. While our token-level deferrals paradigm can apply to all sequence predictors, we will specifically focus on creating a dynamic L2D system with an autoregressive predictor.
In this setting, at the $j^{\text{th}}$ step, the rejector for the $j^{\text{th}}$ token, $r_{j}$, takes in $x$ and $\widehat{y}_{<j}$. It is important to note that $\widehat{y}_{<j}$ can be viewed as a mixture of predictions from the model and the expert. When $r_{j}(x,\widehat{y}_{<j})\geq 0$, the rejector chooses to reject the $j^{\text{th}}$ label prediction from $h$ and receives a prediction $\widehat{y}_{j}^{e}$ from an expert. Otherwise, the model will continue with the prediction from $h$, i.e. $\widehat{y}_{j}^{h}=h(x,\widehat{y}_{<j})$. For a fixed $h$ and a specific instance of $(x,\widehat{y}_{<j},y)$, the optimal $r_{j}$ can be determined by minimizing the following loss function:
$\begin{split}{\cal L}_{j}\left(h,r_{j},x,\widehat{y}_{<j},y\right)=l&\left(y,\widehat{y}_{j}^{h}\right)\mathbbm{1}_{r_{j}\left(x,\widehat{y}_{<j}\right)<0}\\ &+c_{j}(x,\widehat{y}_{<j},y)\mathbbm{1}_{r_{j}\left(x,\widehat{y}_{<j}\right)\geq 0}\end{split}$ (3)
where $l\left(y,\widehat{y}_{j}^{h}\right)$ is the loss incurred from using $\widehat{y}_{j}^{h}$ and $c_{j}(x,\widehat{y}_{<j},y)$ is the cost of querying the expert for the $j^{\text{th}}$ token. $l\left(y,\widehat{y}_{j}^{h}\right)$ can be the mean squared error defined by $(y_{j}-\widehat{y}_{j}^{h})^{2}$ or a 00-$1$ loss between the true $j^{\text{th}}$ token and $\widehat{y}_{j}^{h}$ in text sequences. Generally, $l\left(y,\widehat{y}_{j}^{h}\right)$ can be viewed as token-level feedback to generate an accurate complete sequence. $c_{j}(x,\widehat{y}_{<j},y)$ can be a constant $c_{j}$ to reflect the cost of querying $e$ at the $j^{\text{th}}$ step but it can also be designed as $l\left(y,\widehat{y}_{j}^{e}\right)$ to account for the expert’s errors.
Finally, the predicted $j^{\text{th}}$ label $\widehat{y}_{j}=\widehat{y}_{j}^{h}\mathbbm{1}_{r_{j}\left(x,\widehat{y}_{<j}\right)<0}+\widehat{y}_{j}^{e}\mathbbm{1}_{r_{j}\left(x,\widehat{y}_{<j}\right)\geq 0}$ is then appended to $\widehat{y}_{<j}$ to serve as an input for $(j+1)^{\text{th}}$ token prediction, and this process continues until $L$ labels in the sequence are generated. The total deferral loss for the entire sequence can be characterized as ${\cal L}^{\text{Token}}\left(h,r,x,y\right)=\frac{1}{L}\sum_{j=1}^{L}{\cal L}_{j}\left(h,r_{j},x,\widehat{y}_{<j},y\right)$ where $r=\begin{bmatrix}r_{1}&\dots&r_{L}\end{bmatrix}^{T}$. The ${\cal L}^{\text{Token}}$-risk is denoted by ${\cal R}_{\text{Token}}(h,r)=\mathop{\mathbb{E}}_{(x,y)\sim{\cal P}_{X,Y}}[{\cal L}^{\text{Token}}(h,r,x,y)]$.
3.1 Surrogate Loss
As described in Section 2.2, minimizing the deferral loss directly is computationally difficult, so identifying an optimal rejector requires proposing and optimizing a convex surrogate loss that is Bayes consistent with respect to ${\cal L}^{\text{Token}}$. To this end, we propose the following surrogate loss function: ${\cal L}^{\phi}(h,r,x,y)=\frac{1}{L}\sum_{j=1}^{L}{\cal L}^{\phi}_{j}(h,r_{j},x,\widehat{y}_{<j},y)$ where
$\begin{split}{\cal L}^{\phi}_{j}(h,r_{j},x,\widehat{y}_{<j},y)=l&\left(y,\widehat{y}_{j}^{h}\right)\phi(r_{j}(x,\widehat{y}_{<j}))\\ &+c_{j}(x,\widehat{y}_{<j},y)\phi(-r_{j}(x,\widehat{y}_{<j}))\end{split}$ (4)
and $\phi$ is a convex binary surrogate loss. $\phi$ could, for example, be the logistic loss, i.e. $\phi(z)=\log\left(1+\exp\{-z\}\right)$ or the square loss, i.e. $\phi(z)=(1-z)^{2}$. ${\cal L}^{\phi}$-risk is characterized by ${\cal R}_{\phi}(h,r)=\mathop{\mathbb{E}}_{(x,y)\sim{\cal P}_{X,Y}}[{\cal L}^{\phi}(h,r,x,y)]$.
The optimization of ${\cal L}^{\phi}(h,r,x,y)$ is computationally efficient due its convexity. Additionally, the surrogate loss upper bounds ${\cal L}^{\text{Token}}$ upto a scaling factor, providing alignment between reducing the surrogate loss and reducing the desired loss function. The proof of these properties can be found in Section A.1.
To show Bayes consistency of the surrogate, we relate the excess risk using ${\cal L}^{\text{Token}}$ and the excess risk using ${\cal L}^{\phi}$ in Lemma 3.1. The proof employs a specific type of distribution on ${\cal X}\times\{-1,1\}$ with mass on functions of the conditional expectations of $l\left(y,\widehat{y}_{j}^{h}\right)$ and $c_{j}(x,\widehat{y}_{<j},y)$ to relate the two inequalities - a strategy inspired by Mao et al. [15]. The complete proof can be found in Section A.2.
Lemma 3.1.
Let $0<\bar{c}\leq c_{j}(x,\widehat{y}_{<j},y)\leq\bar{C}<\infty$ for all $j$, $0\leq l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$, and $\phi$ be a binary surrogate loss that satisfies the following inequality for any distribution over ${\cal X}\times\{-1,1\}$ and any measurable function $f$:
${\cal R}_{\text{binary }0-1}(f)-{\cal R}^{*}_{\text{binary }0-1}\leq\Gamma\left({\cal R}_{\text{binary }\phi}(f)-{\cal R}^{*}_{\text{binary }\phi}\right)$
where ${\cal R}_{\text{binary }0-1}$ is the binary 0-1 risk and $\Gamma:\mathbb{R}^{+}\to\mathbb{R}^{+}$ is a non-decreasing concave function. Then, for any measurable $r$ and any $h$, the following inequality holds for any distribution over ${\cal X}\times{\cal Y}$:
${\cal R}_{\text{Token}}\left(r,h\right)-{\cal R}_{\text{Token}}^{*}(h)\leq\tilde{\Gamma}\left({\cal R}_{\phi}\left(r,h\right)-{\cal R}^{*}_{\phi}(h)\right)$
where $\tilde{\Gamma}(z)=\left(\bar{l}+\bar{C}\right)\Gamma\left(\frac{z}{\bar{c}}\right)$, ${\cal R}^{*}_{\text{Token}}(h):=\inf_{r}{\cal R}_{\text{Token}}\left(r,h\right)$, and ${\cal R}^{*}_{\phi}(h):=\inf_{r}{\cal R}_{\phi}\left(r,h\right)$
Taking the limit on both sides of the inequality result of Lemma 3.1 proves consistency of the surrogate loss, formally stated in Theorem 3.2. The proof of Theorem 3.2 hinges on the fact that a classification-calibrated surrogate loss satisfies the condition presented in Lemma 3.1. A loss $\phi$ is considered binary classification calibrated if the minimizer of the pointwise $\phi-$risk or $\mathbbm{E}_{y\sim{\cal P}_{{\cal Y}\mid{\cal X}=x}}[\phi(yf(x))]$ always has the same sign as $\mathbb{P}[Y=1\mid X=x]-\frac{1}{2}$. We defer the full proof of Theorem 3.2 and the formal definition of classification calibrated loss to Section A.3.
Theorem 3.2 (Consistency).
If $0<\bar{c}\leq c_{j}(x,\widehat{y}_{<j},y)\leq\bar{C}<\infty$ for all $j$, $0\leq l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$, and $\phi$ be binary classification calibrated, then, for a fixed $h$, ${\cal L}^{\phi}$ is a Bayes consistent surrogate for ${\cal L}^{\text{Token}}$.
3.2 Generalization Bounds
While Bayes consistency is desirable, the property only holds in the limit of infinite data and does not possess finite sample guarantees. In practice, finding a risk minimizer over the entire class of measurable functions is extremely challenging, particularly because ${\cal P}_{X,Y}$ is usually unknown. With access to a rejector training data set ${\cal D}$ with $n$ samples, we can only learn the optimal $\widehat{r}_{n}\in\mathop{\mathrm{argmin}}\limits_{r\in{\cal F}}\widehat{\cal R}_{\phi}(r)$ from a restricted hypothesis class ${\cal F}$ by minimizing the empirical risk with respect to ${\cal L}^{\phi}$$\widehat{\cal R}_{\phi}(r)=\frac{1}{n}\sum_{(x^{i},y^{i})\in{\cal D}}{\cal L}^{\phi}(h,r,x^{i},y^{i})$. We now present an upper bound for the estimation error to characterize the quality of $\widehat{r}_{n}$ with respect to ${\cal F}$. Please refer to Section A.4 for the full proof.
Theorem 3.3.
Suppose $r\in{\cal F}$ which is composed of $\{{\cal F}_{j}\}_{j=1}^{L}$ such that $r_{j}\in{\cal F}_{j}$, $c_{j}(x)\leq\bar{C}$, and $l(y,\widehat{y}^{h}_{j})\leq\bar{l}$. If $\phi$ is $\rho$-lipschitz continuous, then with probability $1-\delta$, the following upper bound holds for the empirical risk minimizer $\widehat{r}_{n}$ with respect to ${\cal L}_{\phi}$:
$\begin{split}{\cal R}_{\phi}&\left(\widehat{r}_{n}\right)-{\cal R}^{*}_{\phi}\left({\cal F}\right)\\ &\leq\frac{4\rho\sqrt{2\left(\bar{C}^{2}+\bar{l}^{2}\right)}}{L}\sum_{j=1}^{L}\mathfrak{R}_{n}\left({\cal F}_{j}\right)+(\bar{C}+\bar{l})\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}\end{split}$
where ${\cal R}^{*}_{\phi}\left({\cal F}\right)=\inf_{r\in{\cal F}}{\cal R}_{\phi}\left(r\right)$ and $\mathfrak{R}_{n}\left({\cal F}_{j}\right)$ is the Rademacher Complexity of ${\cal F}_{j}$
From Theorem 3.3, the estimation error of using $\widehat{r}_{n}$ with respect to ${\cal L}^{\phi}$ depends on the Rademacher complexity $\mathfrak{R}_{n}\left({\cal F}\right)$ of the hypothesis class. Theorem 3.3 alone does not describe the excess risk with respect to the original token loss ${\cal L}^{\text{Token}}$. Using Lemma 3.1, we express the upper bound of the excess risk in Corollary 3.4, and the proof can be found in Section A.5.
Corollary 3.4.
Suppose $r\in{\cal F}$ which is composed of $\{{\cal F}_{j}\}_{j=1}^{L}$ such that $r_{j}\in{\cal F}_{j}$, $0<\bar{c}\leq c_{j}(x)\leq\bar{C}$, and $l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$. If $\phi$ is a $\rho$-lipschitz continuous binary classification-calibrated function, then with probability $1-\delta$, the following upper bound on the excess risk with respect to ${\cal L}^{\text{Token}}$ holds for the empirical risk minimizer $\widehat{r}_{n}$ with respect to ${\cal L}_{\phi}$ for a fixed $h$:
$\begin{split}{\cal R}_{\text{Token}}\left(\widehat{r}_{n},h\right)-{\cal R}^{*}_{\text{Token}}(h)\leq\tilde{\Gamma}\left({\cal B}({\cal F})+{\cal A}_{\phi}\left({\cal F},h\right)\right)\end{split}$
where $\tilde{\Gamma}$ is a non-decreasing function, ${\cal B}({\cal F})=\frac{4\rho\sqrt{2\left(\bar{C}^{2}+\bar{l}^{2}\right)}}{L}\sum_{j=1}^{L}\mathfrak{R}_{n}\left({\cal F}_{j}\right)+(\bar{C}+\bar{l})\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}$, and ${\cal A}_{\phi}\left({\cal F},h\right)={\cal R}^{*}_{\phi}\left({\cal F},h\right)-{\cal R}^{*}_{\phi}(h)$ is the approximation error.
4 One-time partial deferral
Token-level deferrals are ideal to produce partial deferrals, but they hinge on the expert’s capability to perform next-token prediction instead of generating the entire sequence at once. If the expert is at least capable of completing sequences using predicted subsequences from models, the system can achieve better cost-quality tradeoffs with partial deferrals compared to whole deferrals, albeit constrained by the lack of granularity in deferral, through one-time deferral. In the one-time partial deferral setting, the rejector takes in $x$ as input and identifies a point in sequence - referred to as the deferral point - from which the expert predicts the remaining sequence. When $r(x)=j$, the rejector chooses to reject the prediction of the sequence from the $j^{\text{th}}$ token onward and thereafter defers to the expert. In doing so, the system incurs a loss of $l(y,\widehat{y}_{<j})$ for using the first $j$ tokens from the predictor $h$ and pays a cost of $\tilde{c}_{j}(x,\widehat{y}_{<j},y)$ for using the expert with the partial feedback from $h$. When $r(x)=1$, the entire prediction task is deferred to the expert. If $r(x)=L+1$, the model completes the full prediction without paying any cost for an expert. For this deferral point classification, our objective is to minimize the following loss function:
$\begin{split}{\cal L}^{\text{OneTime}}_{\text{Full}}&\left(h,r,x,y\right)\\ &\sum_{j=1}^{L+1}\left(l(y,\widehat{y}_{<j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)=j}\end{split}$ (5)
where $l(y,\widehat{y}_{<1})=0$
4.1 Influence of large $L$
The one-time deferral paradigm is strongly related to the learning to defer with multiple experts framework, which is evident from the similarities between their loss functions (Equations 2 and 5). Each point of the sequence can be seen as corresponding to a distinct “expert” with its prediction consisting of the predictor output up to that point and to the deferred expert’s prediction for the remainder. However, L2D system accuracy tends to plateau with an increasing number of experts, which naturally has ramifications on the cost-effectiveness of the system [35, 36]. In particular, it suggests that allowing for one-time deferral at each point of the sequence may be over-finely resolved, as deferring on neighboring labels may offer a negligible difference in accuracy.
Towards this end, we wish to allow for flexibility in adjusting the number of “experts”. We achieve this by considering a generalization of one-time deferral where the predictor can defer to the expert on a select set of token positions specified by ${\cal J}$, instead of having all token positions be potential deferral points. For this generalization, we consider the following loss:
$\begin{split}{\cal L}^{\text{OneTime}}&\left(h,r,x,y\right)\\ &=\sum_{j\in{\cal J}}\left(l(y,\widehat{y}_{<j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)=j}\end{split}$ (6)
The ${\cal L}^{\text{OneTime}}$-risk is denoted by ${\cal R}_{\text{OneTime}}(h,r)=\mathop{\mathbbm{E}}_{(x,y)\sim{\cal P}_{X,Y}}[{\cal L}^{\text{OneTime}}(h,r,x,y)]$.
4.2 Surrogate loss
Let ${\boldsymbol{g}}$ consist of $|{\cal J}|$ scoring functions $g_{j}:{\cal X}\to\mathbb{R}$ where $j$ indexes the token positions in a sequence and let $r(x)=\mathop{\mathrm{argmax}}\limits_{j\in{\cal J}}g_{j}(x)$. We propose the following surrogate loss function:
${\cal L}^{\psi}$ $\left(h,r,x,y\right)$
$=\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{<j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\psi({\boldsymbol{g}}(x),j)$
where $\psi({\boldsymbol{g}}(x),j)$ is a consistent surrogate for the multiclass 0-1 loss or $\mathbbm{1}_{r(x)\neq j}$, and $c_{\text{max}}=\max_{j\in{\cal J}}l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)$. Examples of $\psi$ include cross entropy loss, i.e. $\psi_{\text{ce}}({\boldsymbol{g}}(x),j)=-\log\frac{e^{g_{j}(x)}}{\sum_{k\in{\cal J}}e^{g_{k}(x)}}$ or mean absolute error, i.e. $\psi_{\text{mae}}({\boldsymbol{g}}(x),j)=1-\frac{e^{g_{j}(x)}}{\sum_{k\in{\cal J}}e^{g_{k}(x)}}$ [37]. The ${\cal L}^{\psi}$-risk is characterized by ${\cal R}_{\psi}(h,r)=\mathop{\mathbbm{E}}_{(x,y)\sim{\cal P}_{X,Y}}[{\cal L}^{\psi}(h,r,x,y)]$.
Intuitively, for a given token position $j$, $c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)$ is large when the subsequence up till $j$ can be predicted correctly and the cost of deferring on the rest of the sequence is small. When this term is large, we desire $\psi({\boldsymbol{g}}(x),j)$ to be small, incentivizing the model to choose $j$ as the deferral point to minimize the surrogate loss. This surrogate loss encourages the rejector to pick a token position which also has minimal costs for deferring on the rest of the sequence - a goal the ${\cal L}^{\text{OneTime}}$ loss function shares.
Just like ${\cal L}^{\phi}$, ${\cal L}^{\psi}$ is also convex and upper bounds ${\cal L}^{\text{OneTime}}$ up to some scale $\gamma$ (See Section B.2 for proof). Theorem 4.1 shows that ${\cal L}^{\psi}$ achieves Bayes consistency.
Theorem 4.1 (Consistency).
Let $\tilde{c}_{j}(x,\widehat{y}_{<j},y)\leq\bar{C}<\infty$, $l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$ for all $j\in{\cal J}$, and $\psi$ be cross entropy loss $\psi_{\text{ce}}$ or mean absolute error $\psi_{\text{mae}}$. If there exists $i,j\in{\cal J}$ such that $\left|l(y,\widehat{y}^{h}_{i})+\tilde{c}_{i}(x,\widehat{y}_{<i},y)-l(y,\widehat{y}^{h}_{j})-\tilde{c}_{i}(x,\widehat{y}_{<i},y)\right|>\Delta>0$, then, for a fixed $h$, ${\cal L}^{\psi}$ is a consistent surrogate for ${\cal L}^{\text{OneTime}}$.
The proof of the above theorem can be found in Section B.4. If the $l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)$ terms are equal for all $j$, the weights would be 00. By assuming a non-zero cost difference for at least one pair, we are ensuring that at least one $j$ has a non-zero weight. Such an assumption can be satisfied if there is a sufficient separation in predictive capabilities between the predictor and the expert.
4.3 Generalization Bounds
Under the empirical risk minimization framework, the optimal rejector $\widehat{r}_{n}\in\mathop{\mathrm{argmin}}\limits_{r\in{\cal F}}\widehat{\cal R}_{\psi}(r)$ minimizes the empirical risk with respect to ${\cal L}^{\psi}$ or ^$\widehat{\cal R}_{\psi}(r)=\frac{1}{n}\sum_{(x^{i},y^{i})\in{\cal D}}{\cal L}^{\psi}(h,r,x^{i},y^{i})$ from a finite hypothesis class ${\cal F}$. Theorem 4.2 presents generalization bounds for ${\cal L}^{\psi}$ (See Section B.5 for the proof).
Theorem 4.2.
Suppose $r_{n}\in{\cal F}=\{x\to\mathop{\mathrm{argmin}}\limits_{j\in{\cal J}}g_{j}(x):g_{j}\in{\cal G}_{j}\}$ where ${\cal G}_{j}$ consists of functions having a range of $(-M,M)$ with $M\geq 0$, $\tilde{c}_{j}(x,\widehat{y}_{<j},y)\leq\bar{C}$, and $l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$. If $\psi$ is $\rho$-lipschitz continuous with respect to ${\boldsymbol{g}}(x)$ and there exists $u_{\psi}(M)$ such that $\psi\left({\boldsymbol{g}}(x),j\right)\leq u_{\psi}(M)<\infty$ for all $j$ and $r\in{\cal F}$, then with probability $1-\delta$ for a fixed $h$, the following upper bound holds for the empirical risk minimizer $\widehat{r}_{n}$ with respect to ${\cal L}_{\psi}^{\text{Point}}$:
$\begin{split}{\cal R}_{\psi}&\left(\widehat{r}_{n},h\right)-{\cal R}^{*}_{\psi}\left({\cal F},h\right)\\ &\leq 2\sqrt{2}q(L)\rho\sum_{j\in{\cal J}}\mathfrak{R}_{n}\left({\cal G}_{j}\right)+u_{\psi}(M)q(L)\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}\end{split}$
where ${\cal R}^{*}_{\psi}\left({\cal F},h\right)=\inf_{r\in{\cal F}}{\cal R}_{\psi}\left(r,h\right)$, $q(L)=2(|{\cal J}|-1)(\bar{C}+\bar{l})$, and $\mathfrak{R}_{n}\left({\cal G}_{j}\right)$ is the Rademacher Complexity of ${\cal G}_{j}$.
Restricting the hypothesis class to have a range $(-M,M)$ practically involves ReLU activation and softmax clipping to prevent the scores ${\boldsymbol{g}}$ from approaching $\infty$- ∞ or $\infty$. For $\psi_{\text{ce}}$, we can expect $u_{\psi_{\text{ce}}}(M)=-\log\frac{e^{-M}}{e^{-M}+Le^{M}}$ and for $\psi_{\text{mae}}$, it is $u_{\psi_{\text{mae}}}(M)=1-\frac{e^{-M}}{e^{-M}+Le^{M}}$. We can further express the upper bound of the excess risk respect to the original token loss ${\cal L}^{\text{OneTime}}$. The characterization of the upper bound and the proof details can be found in Section B.6.
5 Related Works
Model Cascades [38] share similarities with the learning to defer framework. In cascades, an inference task is progressively handled by a series of models where easier tasks are completed by earlier models in the sequence and complex cases escalate to subsequent models. Similar to confidence-based L2D methods described in Section 2.1, cascading decisions are typically determined by a thresholding deferral rule [39]. Some recent work on applying cascades to large language models [40, 41] has also explored modeling the cascade decision maker. However, model cascades, just like learning to defer methods, ultimately rely on a single model for each prediction. Our work extends cascades by introducing partial deferrals, allowing two models to collaborate on a prediction. Narasimhan et al. [19] also proposes a token-level deferral rule, but their deferral rule compares the predictor’s confidence with expert’s confidence at every token position, which requires many expensive expert queries. Our work learns the expert costs at training time to avoid such expert queries unless the rejector calls for it.
Query Routing methods select a model that provide the highest quality prediction from an ensemble of models. Some query routing algorithms in large language models [42, 43] often require querying either a part or the entire model ensemble to determine the best generation. Some methods train query routing models using specially designed loss functions [44, 45, 46]; these loss functions, however, lack consistency guarantees. Our work not only presents surrogate losses with stronger statistical guarantees but also offers more granular routing decisions.
Many L2D methods draw inspiration from the Cost-Sensitive learning framework. In the example-dependent cost-sensitive setting, each feature $x$ is paired with a cost vector. A cost-sensitive classifier must then learn to identify the index of the cost vector that minimizes the expected cost. Classifiers [47, 48] are often trained with a weighted logistic loss or weighted cross entropy loss function, similar to ${\cal L}^{\text{Token}}$ and ${\cal L}^{\text{OneTime}}$. However, popular methods [49, 50] use losses that are inconsistent with the original cost-sensitive loss.
6 Experiments
In this section, we demonstrate the quality of a model-based approach to partial deferral in two different experimental setups - News Summarization and Traveling Salesman Problem solver. Code will be made available upon acceptance.
6.1 Baselines
In experiments, we wish to highlight two aspects of the proposed methodology. The first is the improved cost-accuracy trade-offs in using partial deferrals over whole-sequence deferrals. The second is the advantages of our specific approach to partial deferrals over confidence-based partial deferrals, i.e. thresholding confidence scores from the predictor.
Towards this end, we conduct two experiments. In the first, we compare our token-level and one-time rejectors with a few whole deferral methods. ChowSum, ChowMean, ChowQuantile are thresholding whole-sequence rejectors of different functions of the negative log probability of the generated sequence. WholeModelEmbed is a model-based whole-sequence rejector trained to predict whether the expert is more accurate than the predictor [40].
For the second experiment, our model-based token-level rejector is compared against two confidence-based deferral methods - TokenwiseSoftmax and TokenwiseEntropy. TokenwiseSoftmax rejector defers prediction when the negative log softmax probability of the greedy token exceeds a threshold, and TokenwiseEntropy makes rejection decisions by thresholding the entropy of the predictive distribution. These methods are similar to the confidence-based methods described in Section 2.1.
In the case of one-time deferral, we similarly consider OneTimeSoftmax and OneTimeEntropy as comparative baselines for the one-time deferral method, which choose the token position based on the lowest softmax probability assigned to the greedy token and the highest entropy, respectively, from a set of tokens, ${\cal J}$. Complete descriptions of all baselines can be found in Appendix C.
6.2 Tasks
We refer the reader to Appendix D for training and hyperparameter details of the rejector models.
Traveling Salesman Problem (TSP): The traveling salesman problem is one of the most widely-known NP-Hard optimization problems which requires finding the shortest route to visit every city exactly once and return to the starting city. While no algorithm can efficiently recover the exact solution to this problem, approximate solutions have proven to be useful practically. These methods include the optimal and time-consuming optimization-based surrogates and efficient, yet less optimal ML-based approaches. We, therefore, use our partial deferral method to facilitate collaborations between both approaches.
In this task, the ML predictor is the PointerNet model of [51] trained on randomly generated TSP graphs with $50$ cities (coordinates are drawn from a standard normal distribution), and the optimization-based expert is a Gurobi solver that takes in the partial solution from the predictor as a constraint. The predicted tours are evaluated by the total distance covered by the tour. Since Gurobi solvers do not solve TSP problems in an autoregressive fashion, producing a single city and completing the whole tour will incur similar costs, making the token-level deferral framework unsuitable in this scenario. This, however, highlights a perfect use-case for one-time deferral. $l(y,\widehat{y}_{<j})$ evaluates the total distance of using the subtour $\widehat{y}_{<j}$ and $\tilde{c}_{j}(x,\widehat{y}_{<j},y)$ is the sum of the total distance of using the expert-generated subtour or $\widehat{y}^{e}_{\geq j}$ and a cost of $\alpha_{j}$ for deferral that increases monotonically with the number of rejected tokens. More details on $\alpha_{j}$ and its effects on performance can be found in Appendix E.
Extreme Summarization (XSUM): In Extreme Summarization [52], the predictor summarizes long articles in a few sentences. In this case, the predictor is a t5-small model and the expert t5-base model with both models finetuned on the XSUM dataset. All the generated summaries are limited to a token length of $20$ and evaluated using ROUGE1 scores.
[…]
6.3 Evaluation
[…]
6.4 Comparison with other deferral methods
Figure 1 and Table 1 show that both model-based token-level deferral and model-based one-time deferral schemes outperform model-based and confidence-based whole deferral schemes with token-level deferral achieving the best performance. Our model-based one-time method also prevails over its counterpart methods and the same holds true for the model-based token-level deferral method. Notably, confidence-based token-level deferral schemes also surpass all other methods in XSUM, demonstrating that token-level deferrals generally offer a superior cost-accuracy tradeoff than whole deferrals. Although model-based one-time deferral exhibits marginal improvement in cost-effectiveness in TSP task, we attribute this to the expert’s lack of ability to predict smaller portions of the sequence at a time. The better cost-accuracy tradeoffs seen in XSUM experiments demonstrate that the advantages of partial deferrals become more prominent with more fine-grained deferrals, which are only possible with access to a suitable expert.
Table 1: Table showing Area under Deferral curve (AUDC) and percentage improvement in AUDC relative to the random rejector for our methods along with the baselines, random, and optimal rejectors. The random rejector makes rejection decisions by flipping a coin with varying probability of landing heads and the optimal rejector achieves the highest possible accuracy regardless of the number of tokens rejected. Numbers in brackets represent the standard deviations measured across the $5$ runs. Low AUDC signifies better cost-accuracy tradeoffs and a positive improvement over a random rejector indicates a shift towards a superlinear tradeoff. The values in bold represent the lowest AUDC and highest percentage improvement for each category (Whole, One-time, Token-level). OneTimeModel scores are reported using the best-performing model trained with ${\cal L}^{\psi}$. See Figure 2 for more details on which performed best.
TSP XSUM
Method AUDC % Improvement AUDC % Improvement
Random 306.12(6.31) 0.00 (0.00) 1433.55 (36.98) 0.00 (0.00)
Optimal 0.00(0.00) 100.00 (0.00) 1212.00 (46.85) 15.48 (1.50)
Whole Deferrals
ChowMean 287.01(7.73) 6.23(2.14) 1430.52 (19.09) 0.13 (3.48)
ChowSum 287.01(7.73) 6.23(2.14) 1431.41 (19.65) 0.06 (3.50)
ChowQuantile 0 301.18(9.40) 1.63(1.18) 1432.28 (25.89) -0.01 (3.89)
ChowQuantile 0.4 287.74(5.91) 6.00(1.16) 1436.10 (24.23) -0.27 (3.80)
ChowQuantile 0.8 288.86(9.35) 5.65(2.12) 1434.48 (19.32) -0.15 (3.40)
ChowQuantile 1.0 295.71(6.12) 3.39(1.42) 1427.27 (18.91) 0.35 (3.50)
WholeModelEmbed 307.88(8.73) -0.58(2.02) 1434.23 (21.57) -0.11 (3.81)
One-time Deferrals
OneTimeModel 270.92(10.67) 11.52(2.41) 1398.39(22.84) 2.35 (4.03)
OneTimeSoftmax 319.79(14.53) -4.44(3.63) 1433.41(26.82) -1.50 (2.13)
OneTimeEntropy 324.23(13.48) -5.93(4.15) 1442.92(24.92) -1.75 (2.38)
Token-level Deferrals
TokenwiseModel - - 1323.92 (25.34) 7.63 (0.93)
TokenwiseSoftmax - - 1354.78 (31.87) 5.48 (0.97)
TokenwiseEntropy - - 1360.41 (31.50) 5.09 (0.63)
Figure 1: Cost-Loss plots for TSP (on the left) and XSUM (on the right) tasks. Deferral curves closer to the lower left corner are more desirable.
6.5 Size of ${\cal J}$ for One time Deferral
We train one-time rejectors using ${\cal L}^{\psi}$ with varying choices of ${\cal J}$ to study the influence of larger sets of deferral point choices on cost-effectiveness. For both XSUM and TSP, ${\cal J}$ represents a grid of deferral points between $1$ and $L+1$ and we vary the granularity of the grid. Figure 2 demonstrates the need for a modified one-time deferral loss function. The model trained with a ${\cal J}$ including all token positions ($|{\cal J}|=L+1$) is not the most optimal choice with the suboptimality exacerbated in the XSUM experiments. One-time rejectors are generally more cost-effective than its corresponding confidence-based rejectors regardless of the size of ${\cal J}$.
Figure 2: Plots of Size of ${\cal J}$ against percentage improvement in AUDC relative to the random baseline for TSP (on the left) and XSUM (on the right). See Appendix F for exact AUDC values with standard deviations.
7 Discussion
We have introduced partial rejections as a way to improve the cost-effectiveness of a learning to defer system and proposed two model-based methods to support granular rejections. This work raises numerous interesting questions. We have currently proposed partial deferral with access to one expert, however, it would be interesting to extend this to the case with multiple experts. It would also be valuable to explore other types of partial deferrals with granularities between token-level and one-time deferrals. A rejector, for example, could select contiguous parts of a sequence to reject. Finally, studying partial rejections for larger outputs like protein structures would also be insightful.
Appendix A Proofs for Section 3
A.1 Proof of Proposition A.1
Proposition A.1.
If $\phi$ is binary classification calibrated and convex, ${\cal L}^{\phi}(h,r,x,y)$ is a convex upper bound of ${\cal L}^{\text{Token}}(h,r,x,y)$ upto some scale $\gamma$
Proof.
It is easy to see that ${\cal L}^{\phi}_{j}(h,r,x,y)$ is convex with respect to $r$ as long as $\phi(\cdot)$ is convex. Since $\phi$ is binary classification calibrated, there is a constant $\gamma$ such that $\gamma\phi(z)\geq\mathbbm{1}_{z\leq 0}$ by Lemma 3 from [53]. Therefore, $\gamma{\cal L}^{\phi}(h,r,x,y)\geq{\cal L}^{\text{Token}}(h,r,x,y)$.
A.2 Proof of Lemma 3.1
Lemma A.2.
Let $0<\bar{c}\leq c_{j}(x,\widehat{y}_{<j},y)\leq\bar{C}<\infty$ for all $j$, $0\leq l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$, and $\phi$ be a binary surrogate loss that satisfies the following inequality for any distribution over ${\cal X}\times\{-1,1\}$ and any measurable function $f$:
${\cal R}_{\text{binary }0-1}(f)-{\cal R}^{*}_{\text{binary }0-1}\leq\Gamma\left({\cal R}_{\text{binary }\phi}(f)-{\cal R}^{*}_{\text{binary }\phi}\right)$
where ${\cal R}_{\text{binary }0-1}$ is the binary 0-1 risk and $\Gamma:\mathbb{R}^{+}\to\mathbb{R}^{+}$ is a non-decreasing concave function. Then, for any measurable $r$ and any $h$, the following inequality holds for any distribution over ${\cal X}\times{\cal Y}$:
${\cal R}_{\text{Token}}\left(r,h\right)-{\cal R}_{\text{Token}}^{*}(h)\leq\tilde{\Gamma}\left({\cal R}_{\phi}\left(r,h\right)-{\cal R}^{*}_{\phi}(h)\right)$
where $\tilde{\Gamma}(z)=\left(\bar{l}+\bar{C}\right)\Gamma\left(\frac{z}{\bar{c}}\right)$, ${\cal R}^{*}_{\text{Token}}(h):=\inf_{r}{\cal R}_{\text{Token}}\left(r,h\right)$, and ${\cal R}^{*}_{\phi}(h):=\inf_{r}{\cal R}_{\phi}\left(r,h\right)$
Proof.
Let $C_{\text{Token}}(r,h,x)=\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[{\cal L}^{\text{Token}}\left(h,r,x,y\right)\right]$ be the pointwise ${\cal L}^{\text{Token}}$- risk for a given $x$ and predictor $h$ and let $C_{\text{Token}}^{*}(h,x)=\inf_{r}C_{\text{Token}}(r,h,x)$ be the optimal pointwise ${\cal L}^{\text{Token}}$- risk for a given $x$ and predictor $h$. Similarly, let $C_{\phi}(r,h,x)=\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[{\cal L}^{\phi}\left(h,r,x,y\right)\right]$ be the pointwise ${\cal L}^{\phi}$- risk for a given $x$ and predictor $h$ and $C_{\phi}^{*}(h,x)=\inf_{r}C_{\phi}(r,h,x)$ be the optimal pointwise ${\cal L}^{\phi}$- risk. For any measurable function $r$, the excess pointwise ${\cal L}^{\phi}$- risk can be characterized by:
$C_{\text{Token}}(r,h,x)-C_{\text{Token}}^{*}(h,x)$
$=\frac{1}{L}\sum_{j=1}^{L}\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[l\left(y,\widehat{y}_{j}^{h}\right)\right]\mathbbm{1}_{r(x,\widehat{y}_{<j})\leq 0}+\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[c_{j}(x,\widehat{y}_{<j},y)\right]\mathbbm{1}_{r(x,\widehat{y}_{<j})>0}$
$\quad-\frac{1}{L}\inf_{\tilde{r}}\left\{\sum_{j=1}^{L}\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[l\left(y,\widehat{y}_{j}^{h}\right)\right]\mathbbm{1}_{\tilde{r}(x,\widehat{y}_{<j})\leq 0}+\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[c_{j}(x,\widehat{y}_{<j},y)\right]\mathbbm{1}_{\tilde{r}(x,\widehat{y}_{<j})>0}\right\}$
Let $A_{j}=\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[l(y,\widehat{y}_{<j})\right]$ and $B_{j}=\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[c_{j}(x,\widehat{y}_{<j},y)\right]$
$C_{\text{Token}}(r,h,x)-C_{\text{Token}}^{*}(h,x)$
$=\frac{1}{L}\sum_{j=1}^{L}A_{j}\mathbbm{1}_{r_{j}(x,\widehat{y}_{<j})\leq 0}+B_{j}\mathbbm{1}_{r_{j}(x,\widehat{y}_{<j})>0}-\frac{1}{L}\inf_{\tilde{r}}\left\{\sum_{j=1}^{L}A_{j}\mathbbm{1}_{\tilde{r}_{j}(x,\widehat{y}_{<j})\leq 0}+B_{j}\mathbbm{1}_{\tilde{r}_{j}(x,\widehat{y}_{<j})>0}\right\}$
$=\frac{1}{L}\sum_{j=1}^{L}\left(A_{j}+B_{j}\right)\left(\frac{A_{j}}{A_{j}+B_{j}}\mathbbm{1}_{r_{j}(x,\widehat{y}_{<j})\leq 0}+\frac{B_{j}}{A_{j}+B_{j}}\mathbbm{1}_{r_{j}(x,\widehat{y}_{<j})>0}\right)$
$\quad-\frac{1}{L}\inf_{\tilde{r}}\left\{\sum_{j=1}^{L}\left(A_{j}+B_{j}\right)\left(\frac{A_{j}}{A_{j}+B_{j}}\mathbbm{1}_{\tilde{r}_{j}(x,\widehat{y}_{<j})\leq 0}+\frac{B_{j}}{A_{j}+B_{j}}\mathbbm{1}_{\tilde{r}_{j}(x,\widehat{y}_{<j})>0}\right)\right\}$
$\leq\frac{\left(\bar{l}+\bar{C}\right)}{L}\sum_{j=1}^{L}\frac{A_{j}}{A_{j}+B_{j}}\mathbbm{1}_{r_{j}(x,\widehat{y}_{<j})\leq 0}+\frac{B_{j}}{A_{j}+B_{j}}\mathbbm{1}_{r_{j}(x,\widehat{y}_{<j})>0}$
$\quad-\frac{\left(\bar{l}+\bar{C}\right)}{L}\inf_{\tilde{r}}\left\{\sum_{j=1}^{L}\frac{A_{j}}{A_{j}+B_{j}}\mathbbm{1}_{\tilde{r}_{j}(x,\widehat{y}_{<j})\leq 0}+\frac{B_{j}}{A_{j}+B_{j}}\mathbbm{1}_{\tilde{r}_{j}(x,\widehat{y}_{<j})>0}\right\}$
Consider a feature space $\tilde{{\cal X}}={\cal X}\times{\cal V}^{*}$ and output space $\tilde{{\cal Y}}=\{-1,1\}$. Let the distribution on $\tilde{X}$ be uniform over $\{\tilde{x}_{1},\cdots,\tilde{x}_{L}\}$ where $\tilde{x}_{j}=(x,\widehat{y}_{<j})$ with $\frac{1}{L}$ mass on each point on the set. Let $P(\tilde{Y}=-1\mid\tilde{X}=\tilde{x}_{j})=\frac{A_{j}}{A_{j}+B_{j}}$ and $f$ be a measurable function such that $f(\tilde{x}_{j})=r_{j}(\tilde{x}_{j})=r_{j}(x,\widehat{y}_{<j})$
Under this distribution,
${\cal R}_{\text{binary }0-1}(r)$ $=\frac{1}{L}\sum_{j=1}^{L}\frac{A_{j}}{A_{j}+B_{j}}\mathbbm{1}_{r_{j}(x,\widehat{y}_{<j})\leq 0}+\frac{B_{j}}{A_{j}+B_{j}}\mathbbm{1}_{r_{j}(x,\widehat{y}_{<j})>0}$
${\cal R}_{\text{binary }0-1}^{*}$ $=\frac{1}{L}\inf_{\tilde{r}}\left\{\sum_{j=1}^{L}\frac{A_{j}}{A_{j}+B_{j}}\mathbbm{1}_{\tilde{r}_{j}(x,\widehat{y}_{<j})\leq 0}+\frac{B_{j}}{A_{j}+B_{j}}\mathbbm{1}_{\tilde{r}_{j}(x,\widehat{y}_{<j})>0}\right\}$
${\cal R}_{\text{binary }\phi}(r)$ $=\frac{1}{L}\sum_{j=1}^{L}\frac{A_{j}}{A_{j}+B_{j}}\phi\left(r_{j}(x,\widehat{y}_{<j})\right)+\frac{B_{j}}{A_{j}+B_{j}}\phi\left(-r_{j}(x,\widehat{y}_{<j})\right)$
${\cal R}_{\text{binary }\phi}^{*}$ $=\frac{1}{L}\inf_{\tilde{r}}\left\{\sum_{j=1}^{L}\frac{A_{j}}{A_{j}+B_{j}}\phi\left(\tilde{r}_{j}(x,\widehat{y}_{<j})\right)+\frac{B_{j}}{A_{j}+B_{j}}\phi\left(-\tilde{r}_{j}(x,\widehat{y}_{<j})\right)\right\}$
From the theorem statement,
$C_{\text{Token}}(r,h,x)-C_{\text{Token}}^{*}(h,x)$
$\leq\frac{\left(\bar{l}+\bar{C}\right)}{L}\sum_{j=1}^{L}\frac{A_{j}}{A_{j}+B_{j}}\mathbbm{1}_{r_{j}(x,\widehat{y}_{<j})\leq 0}+\frac{B_{j}}{A_{j}+B_{j}}\mathbbm{1}_{r_{j}(x,\widehat{y}_{<j})>0}$
$\quad-\frac{\left(\bar{l}+\bar{C}\right)}{L}\inf_{\tilde{r}}\left\{\sum_{j=1}^{L}\frac{A_{j}}{A_{j}+B_{j}}\mathbbm{1}_{\tilde{r}_{j}(x,\widehat{y}_{<j})\leq 0}+\frac{B_{j}}{A_{j}+B_{j}}\mathbbm{1}_{\tilde{r}_{j}(x,\widehat{y}_{<j})>0}\right\}$
$\leq\left(\bar{l}+\bar{C}\right)\Gamma\left(\sum_{j=1}^{L}\frac{A_{j}\phi\left(r_{j}(x,\widehat{y}_{<j})\right)+B_{j}\phi\left(-r_{j}(x,\widehat{y}_{<j})\right)}{L\left(A_{j}+B_{j}\right)}-\inf_{\tilde{r}}\left\{\sum_{j=1}^{L}\frac{A_{j}\phi\left(\tilde{r}_{j}(x,\widehat{y}_{<j})\right)+B_{j}\phi\left(-\tilde{r}_{j}(x,\widehat{y}_{<j})\right)}{L\left(A_{j}+B_{j}\right)}\right\}\right)$
$\leq\left(\bar{l}+\bar{C}\right)\Gamma\left(\sum_{j=1}^{L}\frac{A_{j}\phi\left(r_{j}(x,\widehat{y}_{<j})\right)+B_{j}\phi\left(-r_{j}(x,\widehat{y}_{<j})\right)}{L\bar{c}}-\inf_{\tilde{r}}\left\{\sum_{j=1}^{L}\frac{A_{j}\phi\left(\tilde{r}_{j}(x,\widehat{y}_{<j})\right)+B_{j}\phi\left(-\tilde{r}_{j}(x,\widehat{y}_{<j})\right)}{L\bar{c}}\right\}\right)$
$=\left(\bar{l}+\bar{C}\right)\Gamma\left(\frac{C_{\phi}(r,h,x)-C_{\phi}^{*}(h,x)}{\bar{c}}\right)$
Let $\tilde{\Gamma}(z)=\left(\bar{l}+\bar{C}\right)\Gamma\left(\frac{z}{\bar{c}}\right)$. Then,
${\cal R}_{\text{Token}}\left(r,h\right)-{\cal R}_{\text{Token}}^{*}(h)$ $=\mathop{\mathbb{E}}_{x\sim{\cal P}_{X}}\left[C_{\text{Token}}(r,h,x)-C_{\text{Token}}^{*}(h,x)\right]$
$\leq\mathop{\mathbb{E}}_{x\sim{\cal P}_{X}}\left[\tilde{\Gamma}\left(C_{\phi}(r,h,x)-C_{\phi}^{*}(h,x)\right)\right]$
$\overset{(a)}{\leq}\tilde{\Gamma}\left(\mathop{\mathbb{E}}_{x\sim{\cal P}_{X}}\left[C_{\phi}(r,h,x)-C_{\phi}^{*}(h,x)\right]\right)$
=$\overset{(b)}{=}\tilde{\Gamma}\left({\cal R}_{\phi}\left(r.h\right)-{\cal R}^{*}_{\phi}(h)\right)$
(a) by Jensen’s inequality since $\tilde{\Gamma}$ is concave, (b) since the infimum is taken over all measurable functions ${\cal R}^{*}_{\phi}(h)=\mathop{\mathbb{E}}_{x\sim{\cal P}_{X}}\left[C_{\phi}^{*}(h,x)\right]$
A.3 Proof of Theorem 3.2
Definition A.3 (Classification Calibration).
Let $\eta(x)=\mathbb{P}\left[Y=1\mid X=x\right]$ and ${\cal R}_{\phi}(f)=\mathbb{E}_{x\sim{\cal P}_{X}}\left[C_{\phi}\left(\eta(x),f(x)\right)\right]$ where $C_{\phi}(\eta,t)=\eta\phi(t)+(1-\eta)\phi(-t)$. $\phi$ is considered binary classification calibrated if $C_{\phi}^{-}(\eta)-C^{*}_{\phi}(\eta)>0$ for all $\eta\neq\frac{1}{2}$ where $C_{\phi}^{-}(\eta)=\inf_{t:t\left(\eta-\frac{1}{2}\right)\leq 0}C_{\phi}(\eta,t)$ and $C_{\phi}^{-}(\eta)=\inf_{t\in\mathbbm{R}}C_{\phi}(\eta,t)$
Theorem A.4 (Consistency).
If $0<\bar{c}\leq c_{j}(x,\widehat{y}_{<j},y)\leq\bar{C}<\infty$ for all $j$, $0\leq l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$, and $\phi$ be binary classification calibrated, then, for a fixed $h$, ${\cal L}^{\phi}$ is a Bayes consistent surrogate for ${\cal L}^{\text{Token}}$
Proof.
Since $\phi$ is binary classification calibrated, we know that ${\cal R}_{\text{binary }0-1}(f)-{\cal R}^{*}_{\text{binary }0-1}\leq\psi^{-1}\left({\cal R}_{\text{binary }\phi}(f)-{\cal R}^{*}_{\text{binary }\phi}\right)$ where $\psi$ is a non-decreasing convex function, making $\psi^{-1}$ a non-decreasing concave function [53]. By Lemma 3.1, ${\cal R}_{\text{Token}}\left(r,h\right)-{\cal R}_{\text{Token}}^{*}(h)\leq\tilde{\Gamma}\left({\cal R}_{\phi}\left(r,h\right)-{\cal R}^{*}_{\phi}(h)\right)$ where $\tilde{\Gamma}(z)=\left(1+\bar{C}\right)\psi^{-1}\left(\frac{z}{\bar{c}}\right)$. If $\lim_{n\to\infty}{\cal R}_{\phi}\left(r_{n},h\right)-{\cal R}^{*}_{\phi}(h)=0$, then
$\lim_{n\to\infty}{\cal R}_{\text{Token}}\left(r_{n},h\right)-{\cal R}_{\text{Token}}^{*}(h)$ $\leq\lim_{n\to\infty}\tilde{\psi}\left({\cal R}_{\phi}\left(r_{n},h\right)-{\cal R}^{*}_{\phi}(h)\right)$
=$\overset{(a)}{=}\left(1+\bar{C}\right)\psi\left(\lim_{n\to\infty}\frac{{\cal R}_{\phi}\left(r_{n},h\right)-{\cal R}^{*}_{\phi}(h)}{\bar{c}}\right)$
$=\left(1+\bar{C}\right)\psi(0)$
=$\overset{(b)}{=}0$
(a) by continuity of $\psi$ at 00 and (b) since $\psi(0)=0$ [53]
A.4 Proof of Theorem 3.3
Theorem A.5.
Suppose $r\in{\cal F}$ which is composed of $\{{\cal F}_{j}\}_{j=1}^{L}$ such that $r_{j}\in{\cal F}_{j}$, $c_{j}(x)\leq\bar{C}$, and $l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$. If $\phi$ is $\rho$-lipschitz continuous, then with probability $1-\delta$, the following upper bound holds for the empirical risk minimizer $\widehat{r}_{n}$ with respect to ${\cal L}_{\phi}$ for a fixed $h$:
${\cal R}_{\phi}\left(\widehat{r}_{n},h\right)-{\cal R}^{*}_{\phi}\left({\cal F},h\right)\leq\frac{4\rho\sqrt{2\left(\bar{C}^{2}+\bar{l}^{2}\right)}}{L}\sum_{j=1}^{L}\mathfrak{R}_{n}\left({\cal F}_{j}\right)+(\bar{C}+\bar{l})\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}$
where ${\cal R}^{*}_{\phi}\left({\cal F},h\right)=\inf_{r\in{\cal F}}{\cal R}_{\phi}\left(r,h\right)$ and $\mathfrak{R}_{n}\left({\cal F}_{j}\right)$ is the Rademacher Complexity of ${\cal F}_{j}$
Proof.
Let $G_{n}=\sup_{r\in{\cal F}}\left[{\cal R}_{\phi}\left(r,h\right)-\widehat{\cal R}_{\phi}(r,h)\right]$, $G_{n}^{\prime}=\sup_{r\in{\cal F}}\left[\widehat{\cal R}_{\phi}(r,h)-{\cal R}_{\phi}\left(r,h\right)\right]$, and
=${\cal B}=\{(x,y)\mapsto{\cal L}^{\phi}\left(h,r,x,y\right)\ :r\in{\cal F}\}$.
If $z^{i}=(x^{i},y^{i})$, $G_{n}=g(z^{1},\cdots,z^{n})$. Since $\left|g(z^{1},\cdots,z^{i},\cdots,z^{n})-g(z^{1},\cdots,z^{i^{\prime}},\cdots,z^{n})\right|\leq\frac{\bar{C}+\bar{l}}{n}$, by applying McDiarmid’s inequality
$P\left(G_{n}>\frac{\epsilon}{2}\right)$ $\leq\exp\left\{\frac{-2\left(\frac{\epsilon}{2}-\mathop{\mathbb{E}}\left[G_{n}\right]\right)^{2}}{\sum_{i=1}^{n}\frac{(\bar{C}+\bar{l})^{2}}{n^{2}}}\right\}$
$\leq\exp\left\{\frac{-2n\left(\frac{\epsilon}{2}-2\mathfrak{R}_{n}\left({\cal B}\right)\right)^{2}}{(\bar{C}+\bar{l})^{2}}\right\}$
[…]
Similarly,
$P\left(G_{n}^{\prime}>\frac{\epsilon}{2}\right)$ $\leq\exp\left\{\frac{-2n\left(\frac{\epsilon}{2}-2\mathfrak{R}_{n}\left({\cal B}\right)\right)^{2}}{(\bar{C}+\bar{l})^{2}}\right\}=\frac{\delta}{2}$
Let $\epsilon=4\mathfrak{R}_{n}\left({\cal B}\right)+(\bar{C}+\bar{l})\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}$
$P\left({\cal R}_{\phi}\left(\widehat{r}_{n},h\right)-{\cal R}^{*}_{\phi}\left({\cal F},h\right)>\epsilon\right)$ $\leq P\left(\sup_{r\in{\cal F}}\left|{\cal R}_{\phi}\left(r,h\right)-\widehat{\cal R}_{\phi}(r,h)\right|>\frac{\epsilon}{2}\right)$
$\leq P\left(G_{n}>\frac{\epsilon}{2}\right)+P\left(G_{n}^{\prime}>\frac{\epsilon}{2}\right)$
$\leq\delta$
With at least probability $1-\delta$, ${\cal R}_{\phi}\left(\widehat{r}_{n},h\right)-{\cal R}^{*}_{\phi}\left({\cal F},h\right)\leq 4\mathfrak{R}_{n}\left({\cal B}\right)+(\bar{C}+\bar{l})\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}$.
We can further bound $\mathfrak{R}_{n}\left({\cal B}\right)$ by $\mathfrak{R}_{n}\left({\cal F}_{j}\right)$ if we show that ${\cal L}^{\phi}$ is Lipschitz continuous with respect to $r$
$\left\|{\cal L}^{\phi}(h,r_{1},x,y)-{\cal L}^{\phi}(h,r_{2},x,y)\right\|_{2}^{2}$
$\leq\left\|\frac{1}{L}\sum_{j=1}^{L}l\left(y,\widehat{y}_{j}^{h}\right)\phi(r_{1,j}(x,\widehat{y}_{<j}))-l\left(y,\widehat{y}_{j}^{h}\right)\phi(r_{2,j}(x,\widehat{y}_{<j}))\right\|_{2}^{2}$
$+\left\|\frac{1}{L}\sum_{j=1}^{L}c_{j}(x,\widehat{y}_{<j})\phi(-r_{1,j}(x,\widehat{y}_{<j}))-c_{j}(x,\widehat{y}_{<j})\phi(-r_{2,j}(x,\widehat{y}_{<j}))\right\|_{2}^{2}$
$\leq\frac{\bar{l}^{2}}{L^{2}}\sum_{j=1}^{L}\left\|\phi(r_{1,j}(x,\widehat{y}_{<j}))-\phi(r_{2,j}(x,\widehat{y}_{<j}))\right\|_{2}^{2}+\frac{\bar{C}^{2}}{L^{2}}\sum_{j=1}^{L}\left\|\phi(-r_{1,j}(x,\widehat{y}_{<j}))-\phi(-r_{2,j}(x,\widehat{y}_{<j}))\right\|_{2}^{2}$
$\leq\frac{\rho^{2}(\bar{C}^{2}+\bar{l}^{2})}{L^{2}}\sum_{j=1}^{L}\left\|r_{1,j}(x,\widehat{y}_{<j})-r_{2,j}(x,\widehat{y}_{<j})\right\|_{2}^{2}$
By the extension of the Talagrand’s Contraction Lemma [54],
$\mathfrak{R}_{n}\left({\cal B}\right)\leq\frac{\rho\sqrt{2\left(\bar{C}^{2}+\bar{l}^{2}\right)}}{L}\sum_{j=1}^{L}\mathfrak{R}_{n}\left({\cal F}_{j}\right)$
A.5 Proof of Corollary 3.4
Corollary A.6.
Suppose $r\in{\cal F}$ which is composed of $\{{\cal F}_{j}\}_{j=1}^{L}$ such that $r_{j}\in{\cal F}_{j}$, $0<\bar{c}\leq c_{j}(x)\leq\bar{C}$, and $l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$. If $\phi$ is a $\rho$-lipschitz continuous binary classification-calibrated function, then with probability $1-\delta$, the following upper bound on the excess risk with respect to ${\cal L}^{\text{Token}}$ holds for the empirical risk minimizer $\widehat{r}_{n}$ with respect to ${\cal L}_{\phi}$ for a fixed $h$:
$\begin{split}{\cal R}_{\text{Token}}\left(\widehat{r}_{n},h\right)-{\cal R}^{*}_{\text{Token}}(h)\leq\tilde{\Gamma}\left({\cal B}({\cal F})+{\cal A}_{\phi}\left({\cal F},h\right)\right)\end{split}$
where $\tilde{\Gamma}:\mathbb{R}^{+}\to\mathbb{R}^{+}$, ${\cal B}({\cal F})=\frac{4\rho\sqrt{2\left(\bar{C}^{2}+\bar{l}^{2}\right)}}{L}\sum_{j=1}^{L}\mathfrak{R}_{n}\left({\cal F}_{j}\right)+(\bar{C}+\bar{l})\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}$, and ${\cal A}_{\phi}\left({\cal F},h\right)={\cal R}^{*}_{\phi}\left({\cal F},h\right)-{\cal R}^{*}_{\phi}(h)$ is the approximation error.
Proof.
Excess risk can be decomposed into estimation error and approximation error
${\cal R}_{\phi}\left(r,h\right)-{\cal R}^{*}_{\phi}(h)=\underbrace{{\cal R}_{\phi}\left(r,h\right)-{\cal R}^{*}_{\phi}\left({\cal F},h\right)}_{\text{Estimation Error}}+\underbrace{{\cal R}^{*}_{\phi}\left({\cal F},h\right)-{\cal R}^{*}_{\phi}(h)}_{\text{Approximation Error}}$
From Theorem 3.3, we have an upper bound on the estimation error and so using that, we have the following inequality
${\cal R}_{\phi}\left(r,h\right)-{\cal R}^{*}_{\phi}(h)\leq\frac{4\rho\sqrt{2\left(\bar{C}^{2}+\bar{l}^{2}\right)}}{L}\sum_{j=1}^{L}\mathfrak{R}_{n}\left({\cal F}_{j}\right)+(\bar{C}+\bar{l})\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}+{\cal A}_{\phi}\left({\cal F},h\right)$
Since $\phi$ is binary classification calibrated, we know that
${\cal R}_{\text{binary }0-1}(f)-{\cal R}^{*}_{\text{binary }0-1}\leq\psi^{-1}\left({\cal R}_{\text{binary }\phi}(f)-{\cal R}^{*}_{\text{binary }\phi}\right)$
where $\psi$ is a non-decreasing convex function, making $\psi^{-1}$ a non-decreasing concave function. The proof is complete by using Lemma 3.1
Appendix B Proofs for Section 4
B.1 Proof of Lemma B.1
Lemma B.1.
For any $h$, $r$, $x\in{\cal X}$, $y\in{\cal Y}$, the following equality holds
${\cal L}^{\text{OneTime}}\left(h,r,x,y\right)$ $=\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)\neq j}-(|{\cal J}|-1)c_{\text{max}}+\sum_{j\in{\cal J}}l(y,\widehat{y}_{\leq j})+\sum_{j\in{\cal J}}\tilde{c}_{j}(x,\widehat{y}_{<j},y)$
Proof.
${\cal L}^{\text{OneTime}}\left(h,r,x,y\right)$
$=\sum_{j\in{\cal J}}\left(l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)=j}$
$=\sum_{j\in{\cal J}}\left(l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)-\sum_{j\in{\cal J}}\left(l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)\neq j}$
$=\sum_{j\in{\cal J}}\left(l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)-\sum_{j\in{\cal J}}\left(l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)\neq j}+(|{\cal J}|-1)c_{\text{max}}-(|{\cal J}|-1)c_{\text{max}}$
$=\sum_{j\in{\cal J}}\left(l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)-\sum_{j\in{\cal J}}\left(l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)\neq j}+\sum_{j\in{\cal J}}c_{\text{max}}\mathbbm{1}_{r(x)\neq j}-(|{\cal J}|-1)c_{\text{max}}$
$=\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)\neq j}-(|{\cal J}|-1)c_{\text{max}}+\sum_{j\in{\cal J}}l(y,\widehat{y}_{\leq j})+\sum_{j\in{\cal J}}\tilde{c}_{j}(x,\widehat{y}_{<j},y)$
B.2 Proof of Proposition B.2
Proposition B.2.
${\cal L}^{\psi}(h,r,x,y)$ is a convex upper bound of ${\cal L}^{\text{OneTime}}(h,r,x,y)$ for $\psi\equiv\psi_{\text{ce}}$ and $\psi\equiv\psi_{\text{mae}}$
Proof.
It is easy to see that ${\cal L}^{\psi}(h,r,x,y)$ is convex with respect to $r$ as $\psi_{\text{ce}}(\cdot)$ and $\psi_{\text{mae}}(\cdot)$ are convex.
From Lemma B.1, we know that
${\cal L}^{\text{OneTime}}\left(h,r,x,y\right)$ $=\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)\neq j}-(|{\cal J}|-1)c_{\text{max}}+\sum_{j\in{\cal J}}l(y,\widehat{y}_{\leq j})+\sum_{j\in{\cal J}}\tilde{c}_{j}(x,\widehat{y}_{<j},y)$
$\leq\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)\neq j}+c_{\text{max}}$
If $r(x)\neq j$, then $\psi_{\text{ce}}({\boldsymbol{g}}(x),j)\geq\log 2$ and $\psi_{\text{mae}}({\boldsymbol{g}}(x),j)\geq 0.5$. When $r(x)=j$, $\psi_{\text{ce}}({\boldsymbol{g}}(x),j)\geq 0$ and $\psi_{\text{mae}}({\boldsymbol{g}}(x),j)\geq 0$. Therefore, $\gamma_{\text{ce}}\psi_{\text{ce}}({\boldsymbol{g}}(x),j)\geq\mathbbm{1}_{r\left(x\right)\neq j}$ and $\gamma_{\text{mae}}\psi_{\text{mae}}({\boldsymbol{g}}(x),j)\geq\mathbbm{1}_{r\left(x\right)\neq j}$ where $\gamma_{\text{ce}}=\log 0.5$ and $\gamma_{\text{mae}}=2$. So then,
${\cal L}^{\text{OneTime}}\left(h,r,x,y\right)$ $\leq\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)\neq j}+c_{\text{max}}$
$\leq\gamma\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\psi({\boldsymbol{g}}(x),j)+c_{\text{max}}$
Therefore, $\gamma^{\prime}{\cal L}^{\psi}(h,r,x,y)\geq{\cal L}^{\text{OneTime}}(h,r,x,y)$ as long as $c_{\text{max}}$ is finite
B.3 Proof of Lemma B.3
Lemma B.3.
Let $\tilde{c}_{j}(x,\widehat{y}_{<j},y)\leq\bar{C}<\infty$, $l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$ for all $j\in{\cal J}$, and $\psi$ be a multiclass surrogate loss that satisfies the following inequality for any distribution over ${\cal X}\times{\cal J}$ and any measurable function $f$:
${\cal R}_{\text{multi }0-1}(f)-{\cal R}^{*}_{\text{multi }0-1}\leq\Gamma\left({\cal R}_{\text{multi }\psi}(f)-{\cal R}^{*}_{\text{multi }\psi}\right)$
where ${\cal R}_{\text{multi }0-1}$ is the multiclass 0-1 risk and $\Gamma:\mathbb{R}^{+}\to\mathbb{R}^{+}$ is a non-decreasing concave function. If there exists $i,j\in{\cal J}$ such that $\left|l(y,\widehat{y}^{h}_{i})+\tilde{c}_{i}(x,\widehat{y}_{<i},y)-l(y,\widehat{y}^{h}_{j})-\tilde{c}_{i}(x,\widehat{y}_{<i},y)\right|>\Delta>0$, then for any ${\boldsymbol{g}}$ over all measurable functions and for any $h$, the following inequality holds for any distribution over ${\cal X}\times{\cal Y}$:
${\cal R}_{\text{OneTime}}\left(r,h\right)-{\cal R}_{\text{OneTime}}^{*}(h)\leq\tilde{\Gamma}\left({\cal R}_{\psi}\left(r,h\right)-{\cal R}^{*}_{\psi}(h)\right)$
where $\tilde{\Gamma}(z)=(|{\cal J}|-1)\left(\bar{C}+\bar{l}\right)\Gamma\left(\frac{z}{\Delta}\right)$, ${\cal R}^{*}_{\text{OneTime}}(h)=\inf_{r\in{\cal M}}{\cal R}_{\text{OneTime}}\left(r,h\right)$, and ${\cal R}^{*}_{\psi}(h)=\inf_{r}{\cal R}_{\phi}\left(r,h\right)$
Proof.
Let $C_{\text{OneTime}}(r,h,x)=\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[{\cal L}^{\text{OneTime}}\left(h,r,x,y\right)\right]$ be the pointwise ${\cal L}^{\text{OneTime}}$- risk for a given $x$ and predictor $h$ and let $C_{\text{OneTime}}^{*}(h,x)=\inf_{r}C_{\text{OneTime}}(r,h,x)$ be the optimal pointwise ${\cal L}^{\text{OneTime}}$- risk for a given $x$ and predictor $h$. Similarly, let $C_{\psi}(r,h,x)=\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[{\cal L}^{\text{OneTime}}_{\psi}\left(h,r,x,y\right)\right]$ be the pointwise ${\cal L}^{\psi}$- risk for a given $x$ and predictor $h$ and $C_{\psi}^{*}(h,x)=\inf_{r}C_{\psi}(r,h,x)$ be the optimal pointwise ${\cal L}^{\psi}$- risk for a given $x$ and predictor $h$. For any measurable function ${\boldsymbol{g}}$
$C_{\text{OneTime}}(r,h,x)-C_{\text{OneTime}}^{*}(h,x)$
$=\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[\sum_{j\in{\cal J}}\left(l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)=j}\right]$
$\quad-\inf_{\tilde{r}}\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[\sum_{j\in{\cal J}}\left(l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{\tilde{r}\left(x\right)=j}\right]$
=$\overset{(a)}{=}\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)\neq j}-(|{\cal J}|-1)c_{\text{max}}+\sum_{j\in{\cal J}}l(y,\widehat{y}_{\leq j})+\sum_{j\in{\cal J}}\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right]$
$\quad-\inf_{\tilde{r}}\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{\tilde{r}\left(x\right)\neq j}-(|{\cal J}|-1)c_{\text{max}}+\sum_{j\in{\cal J}}l(y,\widehat{y}_{\leq j})+\sum_{j\in{\cal J}}\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right]$
$=\sum_{j\in{\cal J}}\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right]\mathbbm{1}_{r(x)\neq j}-\inf_{\tilde{r}}\sum_{j\in{\cal J}}\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right]\mathbbm{1}_{\tilde{r}(x)\neq j}$
(a) By Lemma B.1. Let $A_{j}=\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right]$
$C_{\text{OneTime}}(r,h,x)-C_{\text{OneTime}}^{*}(h,x)$
$=\sum_{j\in{\cal J}}A_{j}\mathbbm{1}_{r(x)\neq j}-\inf_{\tilde{r}}\sum_{j\in{\cal J}}A_{j}\mathbbm{1}_{\tilde{r}(x)\neq j}$
$=\sum_{k\in{\cal J}}A_{k}\left(\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\mathbbm{1}_{r(x)\neq j}-\inf_{\tilde{r}}\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\mathbbm{1}_{\tilde{r}(x)\neq j}\right)$
Consider a degenerate distribution on $\tilde{X}$ with a pmf of $1$ on $x$. Let $P(\tilde{Y}=j\mid\tilde{X}=x)=\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}$ and $f$ be a measurable function such that $f(x)=r(x)$. Under this distribution,
${\cal R}_{\text{multi }0-1}(r)=\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\mathbbm{1}_{r(x)\neq j}$ ${\cal R}_{\text{multi }0-1}^{*}=\inf_{\tilde{r}}\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\mathbbm{1}_{\tilde{r}(x)\neq j}$
${\cal R}_{\text{multi }\psi}(r)=\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left({\boldsymbol{g}}(x),j\right)$ ${\cal R}_{\text{multi }\psi}^{*}=\inf_{\tilde{{\boldsymbol{g}}}}\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left({\boldsymbol{g}}(x),j\right)$
From the theorem statement,
$C_{\text{OneTime}}(r,h,x)-C_{\text{OneTime}}^{*}(h,x)$
$=\sum_{k\in{\cal J}}A_{k}\left(\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\mathbbm{1}_{r(x)\neq j}-\inf_{\tilde{r}}\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\mathbbm{1}_{\tilde{r}(x)\neq j}\right)$
$\leq\sum_{k\in{\cal J}}A_{k}\Gamma\left(\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left({\boldsymbol{g}}(x),j\right)-\inf_{\tilde{{\boldsymbol{g}}}}\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left(\tilde{{\boldsymbol{g}}}(x),j\right)\right)$
We know that $\sum_{j\in{\cal J}}A_{j}\leq(|{\cal J}|-1)\left(\bar{C}+\bar{l}\right)$ from:
$\sum_{j\in{\cal J}}A_{j}$ $=\sum_{j\in{\cal J}}\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right]$
$\leq(|{\cal J}-1|)\left(\bar{C}+\bar{l}\right)$
Since $l(y,\widehat{y}^{h}_{i})+\tilde{c}_{i}(x,\widehat{y}_{<i},y)\neq l(y,\widehat{y}^{h}_{j})+\tilde{c}_{i}(x,\widehat{y}_{<i},y)$ for a pair $(i,j)$, we can assume that $\sum_{j\in{\cal J}}A_{j}\geq\Delta$ if $\left|l(y,\widehat{y}^{h}_{i})+\tilde{c}_{i}(x,\widehat{y}_{<i},y)-l(y,\widehat{y}^{h}_{j})-\tilde{c}_{i}(x,\widehat{y}_{<i},y)\right|>\Delta$.
With these bounds,
$C_{\text{OneTime}}(r,h,x)-C_{\text{OneTime}}^{*}(h,x)$
$\leq\sum_{k\in{\cal J}}A_{k}\Gamma\left(\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left({\boldsymbol{g}}(x),j\right)-\inf_{\tilde{{\boldsymbol{g}}}}\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left(\tilde{{\boldsymbol{g}}}(x),j\right)\right)$
$\leq(|{\cal J}|-1)\left(\bar{C}+\bar{l}\right)\Gamma\left(\frac{\sum_{j\in{\cal J}}A_{j}\psi\left({\boldsymbol{g}}(x),j\right)-\inf_{\tilde{{\boldsymbol{g}}}}\sum_{j\in{\cal J}}A_{j}\psi\left(\tilde{{\boldsymbol{g}}}(x),j\right)}{\Delta}\right)$
$=(|{\cal J}|-1)\left(\bar{C}+\bar{l}\right)\Gamma\left(\frac{C_{\psi}(r,h,x)-C_{\psi}^{*}(h,x)}{\Delta}\right)$
Let $\tilde{\Gamma}(z)=(|{\cal J}|-1)\left(\bar{C}+\bar{l}\right)\Gamma\left(\frac{z}{\Delta}\right)$. Then,
${\cal R}_{\text{OneTime}}\left(r,h\right)-{\cal R}_{\text{OneTime}}^{*}(h)$ $=\mathop{\mathbb{E}}_{x\sim{\cal P}_{X}}\left[C_{\text{OneTime}}(r,h,x)-C_{\text{OneTime}}^{*}(h,x)\right]$
$\leq\mathop{\mathbb{E}}_{x\sim{\cal P}_{X}}\left[\tilde{\Gamma}\left(C_{\psi}(r,h,x)-C_{\psi}^{*}(h,x)\right)\right]$
$\overset{(a)}{\leq}\tilde{\Gamma}\left(\mathop{\mathbb{E}}_{x\sim{\cal P}_{X}}\left[C_{\psi}(r,h,x)-C_{\psi}^{*}(h,x)\right]\right)$
=$\overset{(b)}{=}\tilde{\Gamma}\left({\cal R}_{\psi}\left(r,h\right)-{\cal R}^{*}_{\psi}(h)\right)$
(a) by Jensen’s inequality since $\tilde{\Gamma}$ is concave, (b) since the infimum is taken over all measurable functions ${\cal R}^{*}_{\psi}(h)=\mathop{\mathbb{E}}_{x\sim{\cal P}_{X}}\left[C_{\psi}^{*}(h,x)\right]$
B.4 Proof of Theorem 4.1
Theorem B.4 (Consistency).
Let $\tilde{c}_{j}(x,\widehat{y}_{<j},y)\leq\bar{C}<\infty$, $l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$ for all $j\in{\cal J}$, and $\psi$ be cross entropy loss $\psi_{\text{ce}}$ or mean absolute error $\psi_{\text{mae}}$. If there exists $i,j\in{\cal J}$ such that $\left|l(y,\widehat{y}^{h}_{i})+\tilde{c}_{i}(x,\widehat{y}_{<i},y)-l(y,\widehat{y}^{h}_{j})-\tilde{c}_{i}(x,\widehat{y}_{<i},y)\right|>\Delta>0$, then, for a fixed $h$, ${\cal L}^{\psi}$ is a consistent surrogate for ${\cal L}^{\text{OneTime}}$
Proof.
We know that ${\cal R}_{\text{multi }0-1}(f)-{\cal R}^{*}_{\text{multi }0-1}\leq\Gamma_{\text{ce}}\left({\cal R}_{\text{multi }\psi_{\text{ce}}}(f)-{\cal R}^{*}_{\text{multi }\psi_{\text{ce}}}\right)$ from Theorem 3.1 of [55] where $\Gamma_{\text{ce}}^{-1}(z)=\frac{1+z}{2}\log(1+z)+\frac{1-z}{2}\log(1-z)$. Similarly, ${\cal R}_{\text{multi }0-1}(f)-{\cal R}^{*}_{\text{multi }0-1}\leq\Gamma_{\text{mae}}\left({\cal R}_{\text{multi }\psi_{\text{mae}}}(f)-{\cal R}^{*}_{\text{multi }\psi_{\text{mae}}}\right)$ where $\Gamma_{\text{mae}}^{-1}(z)=\frac{z}{L+1}$. Both $\Gamma_{\text{ce}}^{-1}$ and $\Gamma_{\text{mae}}^{-1}$ are non-decreasing convex functions making $\Gamma_{\text{ce}}$ and $\Gamma_{\text{mae}}$ non-decreasing concave functions on $\mathbb{R}^{+}$ domain.
Let $\Gamma\equiv\Gamma_{\text{ce}}$ when $\psi\equiv\psi_{\text{ce}}$ and $\Gamma\equiv\Gamma_{\text{mae}}$ when $\psi\equiv\psi_{\text{mae}}$. By Lemma B.3, ${\cal R}_{\text{OneTime}}\left(r\right)-{\cal R}_{\text{OneTime}}^{*}\leq\tilde{\Gamma}\left({\cal R}_{\psi}\left(r\right)-{\cal R}^{*}_{\psi}\right)$ where $\tilde{\Gamma}(z)=(|{\cal J}|-1)\left(\bar{C}+\bar{l}\right)\Gamma\left(\frac{z}{\Delta}\right)$. If $\lim_{n\to\infty}{\cal R}_{\psi}\left(r_{n}\right)-{\cal R}^{*}_{\psi}=0$, then
$\lim_{n\to\infty}{\cal R}_{\text{Point}}\left(r_{n}\right)-{\cal R}_{\text{Point}}^{*}$ $\leq\lim_{n\to\infty}\tilde{\Gamma}\left({\cal R}_{\psi}\left(r_{n}\right)-{\cal R}^{*}_{\psi}\right)$
=$\overset{(a)}{=}(|{\cal J}|-1)\left(\bar{C}+\bar{l}\right)\Gamma\left(\lim_{n\to\infty}\frac{{\cal R}_{\phi}\left(r_{n}\right)-{\cal R}^{*}_{\phi}}{\Delta}\right)$
$=(|{\cal J}|-1)\left(\bar{C}+\bar{l}\right)\Gamma(0)$
=$\overset{(b)}{=}0$
(a) by continuity of $\Gamma$ at 00 and (b) since $\Gamma_{\text{ce}}(0)=0$ and $\Gamma_{\text{mae}}(0)=0$
B.5 Proof of Theorem 4.2
Theorem B.5.
Suppose $r_{n}\in{\cal F}=\{x\to\mathop{\mathrm{argmin}}\limits_{j\in{\cal J}}g_{j}(x):g_{j}\in{\cal G}_{j}\}$ where ${\cal G}_{j}$ consists of functions having a range of $(-M,M)$ with $M\geq 0$, $\tilde{c}_{j}(x,\widehat{y}_{<j},y)\leq\bar{C}$, and $l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$. If $\psi$ is $\rho$-lipschitz continuous with respect to ${\boldsymbol{g}}(x)$ and there exists $u_{\psi}(M)$ such that $\psi\left({\boldsymbol{g}}(x),j\right)\leq u_{\psi}(M)<\infty$ for all $j\in{\cal J}$ and $r\in{\cal F}$, then with probability $1-\delta$ for a fixed $h$, the following upper bound holds for the empirical risk minimizer $\widehat{r}_{n}$ with respect to ${\cal L}^{\psi}$:
${\cal R}_{\psi}\left(\widehat{r}_{n},h\right)-{\cal R}^{*}_{\psi}\left({\cal F},h\right)\leq 2\sqrt{2}q(L)\rho\sum_{j\in{\cal J}}\mathfrak{R}_{n}\left({\cal G}_{j}\right)+u_{\psi}(M)q(L)\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}$
where ${\cal R}^{*}_{\psi}\left({\cal F},h\right)=\inf_{r\in{\cal F}}{\cal R}_{\psi}\left(r,h\right)$, $q(L)=2(|{\cal J}|-1)(\bar{C}+\bar{l})$, and $\mathfrak{R}_{n}\left({\cal G}_{j}\right)$ is the Rademacher Complexity of ${\cal G}_{j}$.
Proof.
Let $G_{n}=\sup_{r\in{\cal F}}\left[{\cal R}_{\psi}\left(r\right)-\widehat{\cal R}_{\psi}(r)\right]$, $G_{n^{\prime}}=\sup_{r\in{\cal F}}\left[\widehat{\cal R}_{\psi}(r)-{\cal R}_{\psi}\left(r\right)\right]$, and
=${\cal B}=\{(x,y)\mapsto{\cal L}^{\psi}\left(h,r,x,y\right)\ :r\in{\cal F}\}$.
If $z^{i}=(x^{i},y^{i})$, $G_{n}=g(z^{1},\cdots,z^{n})$.
$\left|g(z^{1},\cdots,z^{i},\cdots,z^{n})-g(z^{1},\cdots,z^{i^{\prime}},\cdots,z^{n})\right|$ $\leq\frac{1}{n_{c}}\left|\sup_{r\in{\cal F}}{\cal L}^{\psi}\left(h,r,x^{i},y^{i}\right)-{\cal L}^{\psi}\left(h,r,x^{i^{\prime}},y^{i^{\prime}}\right)\right|$
$\leq\frac{1}{n}\sup_{r\in{\cal F}}\sum_{j\in{\cal J}}\left|\left(c_{\text{max}}^{i}-l(y^{i},\widehat{y}_{\leq j}^{i})-\tilde{c}_{j}(x^{i},\widehat{y}_{<j}^{i},y^{i})\right)\psi({\boldsymbol{g}}(x^{i}),y^{i})\right|$
$\quad+\left|\left(c_{\text{max}}^{i^{\prime}}-l(y^{i^{\prime}},\widehat{y}_{\leq j}^{i^{\prime}})-\tilde{c}_{j}(x^{i^{\prime}},\widehat{y}_{<j}^{i^{\prime}},y^{i^{\prime}})\right)\psi({\boldsymbol{g}}(x^{i^{\prime}}),y^{i^{\prime}})\right|$
$\leq\frac{1}{n}u_{\psi}(M)2(|{\cal J}|-1)(\bar{C}+\bar{l})$
$=\frac{1}{n}u_{\psi}(M)q(L)$
where $q(L)=2(|{\cal J}|-1)(\bar{C}+\bar{l})$. Using this boundedness property, we can apply McDiarmid’s inequality
$P\left(G_{n}>\frac{\epsilon}{2}\right)$ $\leq\exp\left\{\frac{-2\left(\frac{\epsilon}{2}-\mathop{\mathbb{E}}\left[G_{n}\right]\right)^{2}}{\sum_{i=1}^{n_{c}}\frac{u_{\psi}(M)^{2}q(L)^{2}}{n^{2}}}\right\}$
$\leq\exp\left\{\frac{-2n\left(\frac{\epsilon}{2}-2\mathfrak{R}_{n}\left({\cal B}\right)\right)^{2}}{u_{\psi}(M)^{2}q(L)^{2}}\right\}$
[…]
Similarly,
$P\left(G_{n}^{\prime}>\frac{\epsilon}{2}\right)$ $\leq\exp\left\{\frac{-2n\left(\frac{\epsilon}{2}-2\mathfrak{R}_{n}\left({\cal B}\right)\right)^{2}}{u_{\psi}(M)^{2}q(L)^{2}}\right\}=\frac{\delta}{2}$
Then $\epsilon=4\mathfrak{R}_{n}\left({\cal B}\right)+u_{\psi}(M)q(L)\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}$
$P\left({\cal R}_{\psi}\left(\widehat{r}_{n}\right)-{\cal R}^{*}_{\psi}\left({\cal F}\right)>\epsilon\right)$ $\leq P\left(\sup_{r\in{\cal F}}\left|{\cal R}_{\psi}\left(r\right)-\widehat{\cal R}_{\psi}(r)\right|>\frac{\epsilon}{2}\right)$
$\leq P\left(G_{n}>\frac{\epsilon}{2}\right)+P\left(G_{n}^{\prime}>\frac{\epsilon}{2}\right)$
$\leq\delta$
With at least probability $1-\delta$, ${\cal R}_{\psi}\left(\widehat{r}_{n},h\right)-{\cal R}^{*}_{\psi}\left({\cal F},h\right)\leq 4\mathfrak{R}_{n}\left({\cal B}\right)+u_{\psi}(M)q(L)\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}$.
We can further bound $\mathfrak{R}_{n}\left({\cal B}\right)$ by $\mathfrak{R}_{n}\left({\cal F}_{j}\right)$ if we show that ${\cal L}^{\psi}$ is Lipschitz continuous with respect to ${\boldsymbol{g}}$
$\left\|{\cal L}^{\psi}(h,r^{1},x,y)-{\cal L}^{\psi}(h,r^{2},x,y)\right\|_{2}$
$=\left\|\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\psi({\boldsymbol{g}}^{1}(x),j)-\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\psi({\boldsymbol{g}}^{2}(x),j)\right\|_{2}$
$\leq\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\left\|\psi({\boldsymbol{g}}^{1}(x),j)-\psi({\boldsymbol{g}}^{2}(x),j)\right\|_{2}$
=$\overset{(a)}{=}\sum_{j\in{\cal J},j\neq j_{\text{max}}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\left\|\psi({\boldsymbol{g}}^{1}(x),j)-\psi({\boldsymbol{g}}^{2}(x),j)\right\|_{2}$
$\overset{(b)}{\leq}\sum_{j\in{\cal J},j\neq j_{\text{max}}}\left(\bar{C}+\bar{l}\right)\rho\left\|{\boldsymbol{g}}^{1}(x)-{\boldsymbol{g}}^{2}(x)\right\|_{2}$
$=({\cal J}-1)\left(\bar{C}+\bar{l}\right)\rho\left\|{\boldsymbol{g}}^{1}(x)-{\boldsymbol{g}}^{2}(x)\right\|_{2}$
(a) where $j_{\text{max}}=\mathop{\mathrm{argmax}}\limits_{j\in{\cal J}}l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)$, (b) by lipschitz continuity of $\psi$.
By the extension of the Talagrand’s Contraction Lemma [54],
$\mathfrak{R}_{n}\left({\cal B}\right)$ $=\mathop{\mathbb{E}}_{\sigma_{1},...,\sigma_{n},\{z^{i}\}_{i=1}^{n}\sim{\cal P}_{{\cal X}\times{\cal Y}}}\left[\sup_{b\in{\cal B}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}b(z^{i})\right]$
$\leq(|{\cal J}|-1)\left(\bar{C}+\bar{l}\right)\rho\sqrt{2}\mathop{\mathbb{E}}_{\sigma,\{z^{i}\}_{i=1}^{n}\sim{\cal P}_{{\cal X}\times{\cal Y}}}\left[\sup_{{\boldsymbol{g}}\in{\cal G}_{1}\times...\times{\cal G}_{L+1}}\frac{1}{n}\sum_{i=1}^{n}\sum_{j\in{\cal J}}\sigma_{i,j}g_{j}(z^{i})\right]$
$\leq\frac{q(L)\rho}{\sqrt{2}}\sum_{j\in{\cal J}}\mathop{\mathbb{E}}_{\sigma,\{z^{i}\}_{i=1}^{n}\sim{\cal P}_{{\cal X}\times{\cal Y}}}\left[\sup_{g_{j}\in{\cal G}_{j}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i,j}g_{j}(z^{i})\right]$
$\leq\frac{q(L)\rho}{\sqrt{2}}\sum_{j\in{\cal J}}\mathfrak{R}_{n}\left({\cal G}_{j}\right)$
B.6 Proof of Corollary B.6
Corollary B.6.
Suppose $r_{n}\in{\cal F}=\{x\to\mathop{\mathrm{argmin}}\limits_{j\in{\cal J}}g_{j}(x):g_{j}\in{\cal G}_{j}\}$ where ${\cal G}_{j}$ consists of functions having a range of $(-M,M)$ with $M\geq 0$, $\tilde{c}_{j}(x,\widehat{y}_{<j},y)\leq\bar{C}$, and $l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$. If $\psi$ is $\psi_{\text{ce}}$ or $\psi_{\text{mae}}$, then with probability $1-\delta$, for a fixed $h$, the excess risk of with respect to ${\cal L}^{\text{OneTime}}$ holds for the empirical risk minimizer $\widehat{r}_{n}$ with respect to ${\cal L}^{\psi}$
${\cal R}_{\text{OneTime}}\left(\widehat{r}_{n},h\right)-{\cal R}^{*}_{\text{OneTime}}(h)\leq\tilde{\Gamma}\left(2q(L)\rho\sqrt{2}\sum_{j\in{\cal J}}\mathfrak{R}_{n}\left({\cal G}_{j}\right)+u_{\psi}(M)q(L)\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}+{\cal A}_{\psi}\left({\cal F},h\right)\right)$
where $\tilde{\Gamma}:\mathbb{R}^{+}\to\mathbb{R}^{+}$, $q(L)=2(|{\cal J}|-1)(\bar{C}+\bar{l})$, $\rho$ is the lipschitz norm of $\psi$, and ${\cal A}_{\psi}\left({\cal F},h\right)={\cal R}^{*}_{\psi}\left({\cal F},h\right)-{\cal R}^{*}_{\psi}(h)$ is the approximation error.
Proof.
Excess risk can be decomposed into estimation error and approximation error
${\cal R}_{\psi}\left(r,h\right)-{\cal R}^{*}_{\psi}=\underbrace{{\cal R}_{\psi}\left(r,h\right)-{\cal R}^{*}_{\psi}(h)\left({\cal F},h\right)}_{\text{Estimation Error}}+\underbrace{{\cal R}^{*}_{\psi}\left({\cal F},h\right)-{\cal R}^{*}_{\psi}(h)}_{\text{Approximation Error}}$
We know that $\psi_{\text{ce}},\psi_{\text{mae}}$ are upper bounded by $u_{\psi_{\text{ce}}}(M),u_{\psi_{\text{mae}}}(M)$ repsectively. We also know that the lipschitz norm for $\psi_{\text{ce}}$ $\rho_{\text{ce}}$ is $\sqrt{2}$ and the lipschitz norm for $\psi_{\text{mae}}$ $\rho_{\text{mae}}$ is $2$
From Theorem 4.2, we have an upper bound on the estimation error and so using that, we have the following inequality
${\cal R}_{\psi}\left(r,h\right)-{\cal R}^{*}_{\psi}(h)\leq 2q(L)\rho\sqrt{2}\sum_{j\in{\cal J}}\mathfrak{R}_{n}\left({\cal G}_{j}\right)+u_{\psi}(M)q(L)\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}+{\cal A}_{\psi}\left({\cal F},h\right)$
${\cal R}_{\text{multi }0-1}(f)-{\cal R}^{*}_{\text{multi }0-1}\leq\Gamma_{\text{ce}}\left({\cal R}_{\text{multi }\psi_{\text{ce}}}(f)-{\cal R}^{*}_{\text{multi }\psi_{\text{ce}}}\right)$ from Theorem 3.1 of [55] where $\Gamma_{\text{ce}}^{-1}(z)=\frac{1+z}{2}\log(1+z)+\frac{1-z}{2}\log(1-z)$. Similarly, ${\cal R}_{\text{multi }0-1}(f)-{\cal R}^{*}_{\text{multi }0-1}\leq\Gamma_{\text{mae}}\left({\cal R}_{\text{multi }\psi_{\text{mae}}}(f)-{\cal R}^{*}_{\text{multi }\psi_{\text{mae}}}\right)$ where $\Gamma_{\text{mae}}^{-1}(z)=\frac{z}{L+1}$. Both $\Gamma_{\text{ce}}^{-1}$ and $\Gamma_{\text{mae}}^{-1}$ are non-decreasing convex functions making $\Gamma_{\text{ce}}$ and $\Gamma_{\text{mae}}$ non-decreasing concave functions on $\mathbb{R}^{+}$ domain.
The proof is complete by using Lemma B.3
Appendix C Baseline Details
Let $p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)$ be the predictive distribution of the $j^{\text{th}}$ token conditioned on the leftward context $\widehat{y}_{<j}$ and input $x$. The predictive entropy would then be $H(\widehat{Y}_{j}\mid\widehat{y}_{<j},x)=-\sum_{v\in{\cal V}}p_{h}(v\mid\widehat{y}_{<j},x)\log p_{h}(v\mid\widehat{y}_{<j},x)$. For a given threshold $\tau$, the whole deferral decision rule is determined by evaluating $r(x)>\tau$. The baselines offer various ways of modeling $r$.
  • ChowSum - $r_{\text{ChowSum}}(x)=-\sum_{j=1}^{L}\log p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)$
  • ChowMean - $r_{\text{ChowMean}}(x)=-\frac{1}{L}\sum_{j=1}^{L}\log p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)$
  • ChowQuantile at the level $\alpha$ $r_{\text{ChowQuantile}}^{\alpha}(x)=\text{Quantile}_{\alpha}\left(\left\{-\log p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)\right\}_{j=1}^{L}\right)$
  • WholeModelEmbed - $r(x)$ is trained using the surrogate of the modified One-time deferral loss or ${\cal L}^{\psi}_{\text{mod}}$ where ${\cal J}=\{1,L+1\}$. This is quite similar to the approach proposed by Gupta et al. [40] except we train the classifier with a cost-sensitive loss
Similarly for token-wise deferral, if $r_{j}(x,\widehat{y}_{<j})$ exceeds threshold $\tau$, the expert is called to predict the next token. $r_{j}$ can be modeling in the following ways:
  • TokenwiseSoftmax: $r_{j}^{\text{softmax}}(x,\widehat{y}_{<j})=-\log p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)$
  • TokenwiseEntropy: $r_{j}^{\text{entropy}}(x,\widehat{y}_{<j})=H(\widehat{Y}_{j}\mid\widehat{y}_{<j},x)$
Since $r(x)=\mathop{\mathrm{argmax}}\limits_{j\in{\cal J}}g_{j}(x)$ for one-time deferral, we consider the following the scoring functions $g_{j}$:
  • OneTimeSoftmax: $g_{j}^{\text{softmax}}(x)=-\log p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)$
  • OneTimeEntropy: $g_{j}^{\text{entropy}}(x,\widehat{y}_{<j})=H(\widehat{Y}_{j}\mid\widehat{y}_{<j},x)$
Appendix D Training Details and Hyperparameters
All the models were implemented using PyTorch and all the models were trained on an Nvidia A40 GPU. They were trained using Early stopping and Table 2 contains all hyperparameter details.
Task/Hyperparameters TSP-OneTime XSUM-Tokenlevel XSUM-OneTime
Learning rate 5e-4 1e-3 5e-4
Early Stopping (patience, delta) 20, 1e-4 7, 1e-4 20, 1e-4
Architecture MLP LSTM MLP
Hidden layers 1 3 1
Hidden units 8 128 8
Dropout 0.2 0.4 0.2
Epochs 200 100 200
Training Samples 2400 800 2400
Gradient Clipping Norm 1.0 1.0 1.0
Weight Decay 0.005 0.001 0.005
Table 2: Hyperparameter settings
TSP: Using the RL4CO library [56], TSP graphs with $50$ nodes were generated by sampling the node coordinates from a standard normal distribution. These samples were fed into a PointerNet (predictor) trained with $10000$ samples. The node embeddings, graph context, and log scores from the PointerNet served as inputs for the One-Time rejector. This rejector only took a maximum of $10$ minutes to train.
XSUM: During training and inference, Tokenlevel rejector was only exposed to samples with summaries of less than $15$ words. The model was trained without teacher forcing. It took about $5$ hours to train. The token level rejector $r_{1}$ recieves the encoder hidden states from the predictor as hidden states and the decoder hidden states as input. This continues in an autoregressive fashion. We defer the reader to Appendix G for ablation studies on various training strategies for the tokenlevel rejector.
The OneTime rejector was trained on a dataset with summaries containing upto $20$ words and it took a couple of minutes to train after a dataset of all the varying mixtures of predictions was created (this took a couple of hours to make). This rejector takes the encoder hidden states, decoder hidden states, and log scores from the predictor as input.
Appendix E Effects of $\alpha_{j}$
In experiments, we determine the cost of deferring a complete prediction and call it $\alpha_{1}$. Then, we set $\alpha_{j}=\frac{L-j+1}{L}\alpha_{1}$, making the cost of deferring a token $\frac{\alpha_{1}}{L}$ or $\alpha_{L}$. We vary $\alpha_{1}$ and Figure 3 shows how the costs affect cost-accuracy tradeoffs. For both experiments, extreme values for cost often result in poor cost-accuracy tradeoff. For TSP, the results are best when $\alpha_{L}=0.01$ and for XSUM, it is either $\alpha_{L}=0.01$ or $\alpha_{L}=0.015$.
As a general rule of thumb, we propose that $\alpha_{1}$ should be determined by the median difference between the expert accuracy and predictor accuracy on the train set.
Figure 3: Plots of Size of ${\cal J}$ against AUDC and percentage improvement in AUDC relative to the random baseline for TSP (on the left) and XSUM (on the right) for models trained with different $\alpha_{1}$ value. The costs displayed in the legend refer to the deferral cost per token or $\alpha_{L}$
Appendix F AUDC and Percent Improvement for Size of ${\cal J}$ experiment
In support of Figure 2, Table 4 and Table 3 report the actual values of AUDC and percent improvement in AUDC over the random baseline for each size of ${\cal J}$ considered for both TSP and XSUM.
Size of ${\cal J}$ OneTimeModel (Ours) OneTimeSoftmax OneTimeEntropy
AUDC % Improvement AUDC % Improvement AUDC % Improvement
5 287.37(7.67) 6.13(1.68) 322.08(18.24) -5.19(4.99) 316.76(19.90) -3.43(5.30)
10 277.44(10.59) 9.38(2.54) 318.00(12.62) -3.86(2.95) 317.84(22.77) -3.78(6.42)
15 277.62(9.15) 9.32(1.85) 320.45(6.89) -4.68(1.32) 317.87(12.57) -3.83(3.22)
20 273.34(11.13) 10.74(2.34) 321.35(19.06) -4.96(5.43) 320.54(22.24) -4.70(6.68)
25 273.31(12.20) 10.73(3.11) 318.49(12.65) -4.03(3.18) 321.32(13.95) -4.97(4.10)
30 271.10(9.77) 11.45(2.16) 319.61(10.82) -4.39(1.99) 321.67(14.16) -5.05(3.11)
35 273.30(6.76) 10.72(1.69) 316.40(15.66) -3.36(4.55) 320.51(20.33) -4.70(6.20)
40 272.57(10.08) 10.97(2.53) 318.99(14.07) -4.18(3.49) 320.91(15.41) -4.82(4.16)
45 270.92(10.67) 11.52(2.41) 318.52(14.10) -4.04(3.71) 322.02(14.99) -5.19(4.31)
51 272.64(9.11) 10.95(2.06) 318.96(14.78) -4.17(3.66) 322.91(14.05) -5.50(4.39)
Table 3: AUDC scores and Percent improvement over the random baseline in TSP of various sizes of ${\cal J}$ with standard deviations in brackets
Size of ${\cal J}$ OneTimeModel (Ours) OneTimeSoftmax OneTimeEntropy
AUDC % Improvement AUDC % Improvement AUDC % Improvement
3 1412.16(26.70) 1.54(0.88) 1444.68(16.16) -0.74(0.77) 1441.52(18.29) -0.52(0.61)
4 1398.39(25.51) 2.50(0.80) 1448.43(23.88) -1.00(0.91) 1447.35(24.59) -0.92(1.02)
5 1407.01(19.33) 1.89(0.43) 1457.76(23.95) -1.65(1.11) 1450.36(22.32) -1.13(1.04)
7 1406.88(19.64) 1.90(0.79) 1448.54(20.05) -1.01(0.97) 1439.58(16.68) -0.39(1.20)
10 1407.96(19.47) 1.83(0.34) 1452.68(15.51) -1.31(1.44) 1457.43(15.44) -1.63(1.29)
13 1408.87(23.47) 1.76(0.71) 1445.25(12.84) -0.80(1.78) 1450.64(14.74) -1.17(1.72)
15 1408.94(16.75) 1.75(0.68) 1453.11(6.66) -1.34(1.46) 1460.17(14.20) -1.83(1.43)
17 1418.97(16.00) 1.05(0.59) 1450.95(11.65) -1.19(1.63) 1457.15(10.29) -1.62(1.52)
21 1417.50(26.37) 1.17(0.68) 1452.79(15.86) -1.31(1.30) 1447.35(12.72) -0.93(1.20)
Table 4: AUDC scores and Percent improvement over the random baseline in XSUM of various sizes of ${\cal J}$ with standard deviations in brackets
Appendix G Token-level Rejector Ablation
${\cal L}^{\text{Token}}$ bears resemblance to a weighted multi-label classification loss function. Specifically, the token-level rejector follows a classifier chain approach [57]. Classifier chains is a multi-label classification (MLC) approach that involves training $L$ binary classifiers in a pre-specified order with the $j^{\text{th}}$ binary classifier taking the feature vector and the previous $j-1$ labels as input, forming a chain of predictors. At inference time, the test features are fed into the chain such that the previous label predictions are appended to the features for the next label classifier. Our token-level rejector $r_{j}$ implicitly makes these chained predictions by taking in $\widehat{y}_{<j}$ which is informed by the previous $(j-1)$ rejectors’ rejection decision.
With this comparison established, model architectures and training techniques from classifier chain literature can be utilized. Recurrent classifier chains [58] have emerged as a popular choice for MLC as they efficiently maximize accuracy without requiring the large number of trainable parameters typically needed for $K$ separate binary classifiers.
Since the inputs of recurrent classifier chains at training time are typically ground truth labels, training token-level rejectors would similarly require using the correct rejection decision, $\mathbbm{1}_{l\left(y,\widehat{y}_{j}^{h}\right)>c_{j}(x,\widehat{y}_{<j},y)}$, to determine the $j^{\text{th}}$ token prediction for the next label rejector. Following this strategy, also called “teacher forcing” [59], can often cause a distribution shift, as label predictions will be used at test time instead of the correct label, resulting in errors propagating through the chain. Scheduled sampling [27] offers a middle ground by labeling the training data with the correct rejection decision with probability $p$, and otherwise using the predicted one. This probability $p$ often decays at a linear or exponential rate over epochs, gradually reducing the model’s dependence on ground-truth rejection rules.
We use both Multi Layer Perceptrons (MLP) and Long Short-term memory (LSTM) models to perform token-level deferrals on XSUM data. All models are trained with teacher forcing, without teacher forcing, or with scheduled sampling. The Scheduled sampler decayed the teacher forcing probability exponentially by $0.95$ with every epoch and the decay process stopped when the teacher forcing probability was $0.5$. The sampler has $5$ warmup epochs where the teacher forcing probability was $1.0$ before the decay started. The coin flip of whether or not to teacher force occurred on a token-level, as opposed to a batch-level. Figure 4 shows the superiority of recurrent models like LSTMs over MLPs with teacher forcing adversely affecting their performance. The selected model is an LSTM trained without teacher forcing. We present the AUDC scores for each of the models in Table 5.
Figure 4: Cost-Loss plots for XSUM for various model architectures and training strategies.
Method AUC Score % Improvement
MLP No Teacher Forcing 1375.48 (34.57) 4.07 (1.01)
MLP Teacher Forcing 1387.49 (25.96) 3.22 (1.25)
MLP Scheduled Sampling 1377.01 (31.92) 3.96 (1.20)
LSTM No Teacher Forcing 1323.92 (25.34) 7.66 (0.96)
LSTM Teacher Forcing 1340.43 (20.24) 6.49 (1.24)
LSTM Scheduled Sampling 1324.83 (27.04) 7.60 (0.98)
TokenwiseSoftmax 1354.78 (31.87) 5.51 (0.94)
TokenwiseEntropy 1360.41 (31.50) 5.12 (0.62)
Table 5: AUC Scores for various Token-level Rejector architecturs and training